### v1

# Demystifying Invariant Effectiveness for Securing Smart Contracts

Abstract.
Smart contract transactions associated with security attacks often exhibit distinct behavioral patterns compared with historical benign transactions before the attacking events. While many runtime monitoring and guarding mechanisms have been proposed to validate invariants and stop anomalous transactions on the fly, the empirical effectiveness of the invariants used remains largely unexplored. In this paper, we studied $23$ prevalent invariants of $8$ categories, which are either deployed in high-profile protocols or endorsed by leading auditing firms and security experts. Using these well-established invariants as templates, we developed a tool Trace2Inv which dynamically generates new invariants customized for a given contract based on its historical transaction data. We evaluated Trace2Inv on $42$ smart contracts that fell victim to $27$ distinct exploits on the Ethereum blockchain. Our findings reveal that the most effective invariant guard alone can successfully block $18$ of the $27$ identified exploits with minimal gas overhead. Our analysis also shows that most of the invariants remain effective even when the experienced attackers attempt to bypass them. Additionally, we studied the possibility of combining multiple invariant guards, resulting in blocking up to $23$ of the $27$ benchmark exploits and achieving false positive rates as low as $0.32\%$. Trace2Inv outperforms current state-of-the-art works on smart contract invariant mining and transaction attack detection in terms of both practicality and accuracy. Though Trace2Inv is not primarily designed for transaction attack detection, it surprisingly found two previously unreported exploit transactions, earlier than any reported exploit transactions against the same victim contracts.
copyright: nonecopyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXbooktitle: Companion Proceedings of the 32nd ACM Symposium on the Foundations of Software Engineering (FSE ’24), July 15–19, 2024, Porto de Galinhas, Brazilprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Blockchain technology has paved the way for decentralized, resilient, and programmable ledgers on a global scale. One of its most impactful applications is smart contracts, which can be deployed onto a blockchain. These smart contracts allow developers to encode intricate transaction rules that govern the ledger. This innovation has made both blockchains and smart contracts essential infrastructure for decentralized financial services, commonly known as DeFi. As of Sept 25, 2023, the Total Value Locked (TVL) in $2,933$ DeFi protocols has reached an impressive $48.58$ billion (def, 2023a).
However, the landscape is not without its challenges. Security attacks pose a significant threat to the security of smart contracts. Attackers can exploit various vulnerabilities by sending malicious transactions, potentially leading to the theft of millions of dollars from users. As of Sept 25, 2023, the financial losses attributed to security attacks on DeFi protocols exceeded $5.53$ billion USD (def, 2023b).
One key observation is that transactions initiated by attackers often display abnormal behaviors when compared to standard transactions from regular DeFi contract users. These malicious transactions may exploit control flows in corner cases, use abnormally large values to trigger overflows, or manipulate a large volume of digital assets to distort the market in DeFi contracts. In fact, industry experts have been actively monitoring abnormal digital asset movements on-chain to report malicious activities. For example, Forta Network (Network, 2023) deploys monitoring bots to detect on-chain security-related events in real-time. Driven by this observation, smart contract developers have proposed deploying runtime checks to detect transactions leading to abnormal behaviors to neutralize malicious attacks. These checks involve enforcing various runtime invariants, such as restricting the maximum number of digital asset deposits or withdrawals in a contract to prevent market manipulation. Another example is to limit the interaction of other contracts to prevent attackers from crafting sophisticated attack strategies. However, these mechanisms are often manually designed and tailored for specific contracts. This raises questions about their effectiveness across different types of contracts and whether they maintain an acceptable false positive rate without hindering normal user activities.
Smart Contract Invariant Study: This paper presents the first comprehensive, quantitative analysis focused on the utilization of dynamically inferred invariants to enhance smart contract security. We examine $23$ invariant templates, which are advocated by leading auditing firms, academic research, and DeFi protocol developers. Our findings indicate that dynamically inferred dynamic invariants serve as effective mechanisms for thwarting security breaches. When appropriately configured, these invariants can neutralize over 74.1% of malicious attacks while maintaining a false positive rate of less than 0.32%.
Trace2Inv: To facilitate this study, we have developed Trace2Inv, a scalable and extensible invariant synthesis framework. Trace2Inv is designed to automatically derive invariants from transaction traces through the use of trace and dynamic taint analysis. Trace2Inv leverages the main feature of public blockchains, transparent databases of transactions histories containing well-organized transaction execution data. Then, for each invariant template under consideration, Trace2Inv employs a specialized inference algorithm to dynamically generate the corresponding invariant based on historical transaction data.
Experimental Results: We evaluate Trace2Inv on a benchmark set of $42$ smart contracts that have previously fallen victim to security attacks. Our results show that properly constructed invariants are effective in neutralizing security threats in $39$ out of the $42$ benchmark contracts.
In the course of our study, we categorized the $23$ invariant templates into eight distinct groups based on their underlying design principles: access control, time lock, gas control, re-entrancy, oracle, storage, money flow, and data flow. Subsequently, we conducted a series of in-depth analyses to compare the efficacy of invariants within each group. We also manually scrutinized the transactions flagged by each invariant template, leading to several key findings:
  • Finding 1: Certain invariant outperform others in terms of effectiveness. Within each invariant group, we identified at least one invariant template that is quantitatively superior, neutralizing a greater number of attacks while generating fewer false positives. See Section 6.1.
  • Finding 2: Invariants remain effective even when attackers are aware of them in the majority of cases. A common concern regarding runtime invariants is their potential vulnerability to informed attackers. Our study reveals that selected invariants in the access control, time lock, gas control, money flow, and data flow groups often directly counter critical elements of attack strategies, such as flash loans and transaction atomicity. These invariants not only neutralize the malicious transactions but also render the attack strategies unfeasible or non-profitable in 84.21% of cases. See Section 6.2.
  • Finding 3: Normal users can possibly circumvent invariant guards, thereby mitigating the impact on user experience. For example, in the case of data flow and money flow invariants, a user can divide a large transaction into smaller segments to bypass the invariant guard in 80% false positive instances. See Section 6.2.
  • Finding 4: Combined invariants, formed through disjunction or conjunction, offer enhanced security coverage with lower false positive rate. Different groups of invariants address different attack scenarios. A combined invariant formed through conjunction can cover more attack vectors, while one formed through disjunction may reduce the false positive rate, as malicious attacks often exhibit multiple abnormal behaviors. See Section 6.3.
Contributions: This paper presents the following contributions:
  • Invariant Inference: This paper conducts an extensive study of $23$ invariant templates, categorized into $8$ distinct groups. Additionally, we introduce innovative techniques for the effective inference of invariants across all studied templates from transaction history.
  • Trace2Inv: This paper presents the design and implementation of Trace2Inv, a specialized tool for smart contract trace analysis that is capable of inferring the invariants under study from transaction history.
  • Experimental Results: This paper presents the first systematic and quantitative evaluation of the effectiveness of runtime invariants on $42$ victim contracts in $27$ real-world exploits with high financial losses.
  • Invariant Study Findings: Our research uncovers a series of critical insights that will inform the future application and development of dynamic invariants.
2. Background
Blockchain is a distributed, immutable ledger technology that records transactions across multiple nodes in a network. It employs cryptographic techniques to ensure data integrity and consensus algorithms to maintain final consistency across all participating nodes. Smart contracts are self-executing contracts with the terms of the agreement directly written into code. Deployed on blockchains, they are immutable and transparent, enabling trustless transactions without the need for intermediaries. Invariant guards (also called circuit breakers) are runtime checks around contract invariant conditions that shall always hold during contract exeuction, aiming to secure smart contracts on the fly.
Ethereum Virtual Machine (EVM) is the runtime environment for smart contract execution on Ethereum. It is a Turing-complete virtual machine that interprets and executes the bytecode compiled from contracts programmed in high-level language like Solidity. Gas is a unit of transaction fee on Ethereum, used to quantify the computational efforts for the execution of EVM operations. It is paid in Ether, the native cryptocurrency of Ethereum. Externally Owned Account (EOA) and contract account are two types of accounts on Ethereum. EOAs are owned by normal users only who have the right to send transactions to blockchains, while contract accounts are controlled by the code deployed at a certain blockchain address and its code will be executed when the contract function is invoked. ERC20 is a standard interface for fungible tokens on Ethereum. Almost all valuable tokens on Ethereum are ERC20 tokens.
Common smart contract vulnerabilities include integer overflow/underflow (Blockchain-Projects, 2020), reentrancy (Siegel, 2016), dangerous delegatecall (Santiago, 2017), etc. DeFi vulnerabilities are smart contract vulnerabilities that are specific to DeFi applications. They are more subtle to detect and attacks usually involve more sophisticated steps (Zhou et al., 2023). DeFi protocols are major targets for smart contract attacks, which experienced \$5.53 billion loss out of the overall \$6.94 billion in recent blockchain incidents (def, 2023b).
3. Motivating Example
In this section, we present a motivating example to illustrate how an exploit transaction behave differently from other benign transactions in histories and how dynamically inferred transactions can neutralize the exploit transaction to enhance smart contract security.
Exploit Transaction: Harvest Finance is a Decentralized Finance (DeFi) protocol deployed on Ethereum to manage and auto-invest stable coins for users. On October 26, 2020, USDC and USDT vaults of the Harvest Finance were exploited, causing a financial loss of about USD \$$33.8$ million. Harvest Finance internally uses the market data of Curve, another DeFi stable coin trading protocol, to determine the market prices of USDC and USDT. In the exploit transaction, the attacker distorts the market of Curve to cause the Harvest Finance to make sub-optimal investment decisions.
Specifically, the attacker in the exploit transaction first borrows a large amount of digital assets and uses the borrowed asset to buy USDC in Curve to inflate the USDC price in Curve. Then it deposits $49.98$M USDC into Harvest vault contract, which increases its USDC balance from $72.83$M to $122.51$M. Due to the manipulated oracle price of USDC, Harvest vault contract erroneously mints the attacker an inflated $51.46$M fUSDC, which increases the total fUSDC supply from $127.58$M to $179.04$M. The attacker then restores the USDC by selling the USDC in Curve, and redeems all its fUSDC tokens for $50.30$M USDC, yielding a $32$k surplus compared to the initial deposit. This redemption decreases the Harvest vault’s USDC balance from $122.81$M to $72.51$M and restores the total fUSDC supply to its original value of $127.58$M. Remarkably, this identical attack vector is executed three times within one exploit transaction, consuming an unusually high gas count of $9,895,111$, narrowly within the gas limit of $12,065,986$ at the time.
Abnormal Behaviors: We identify four distinct dimensions of abnormal behavior: a high frequency of user interactions with the Harvest vault contract, an exceptionally large volume of token flow, abrupt fluctuations in the total supply of fUSDC tokens, and remarkably high gas consumption. To better understand the abnormality of the exploit transaction, we collect and analyze all transaction history of the Harvest vault contract up to the point of the exploit, as illustrated in Figure 1.
Figure 1. Statistics of Transactions on Harvest USDC Vault Contract.
As shown in Figure 1, the last data point in each sub-figure, representing the exploit transaction, consistently emerges as an outlier. Specifically, we have observed that the exploit transaction is the first in the contract’s history to: (1) invoke the withdraw function from a contract rather than from a user address, (2) call both deposit and withdraw functions $3$ times within one transaction, (3) consume more gas than any previous transaction, (4) withdraw more USDC from the protocol than any other transaction, (5) elevate the total supply of fUSDC tokens to an all-time high.
Apply Inferred Invariants: The multi-dimensional abnormalities observed in the exploit transaction highlight a stark departure from typical transactional behaviors. This divergence suggests the feasibility of crafting and applying runtime invariants that are capable of flagging and blocking transactions exhibiting such anomalous characteristics. For example, suppose we inferred and enforced an invariant stating that the withdraw function may only be invoked by an Externally Owned Account (EOA), the exploit transaction could be blocked. This is because the exploit relies on a contract to execute complex logic designed to extract funds from the vault. Likewise, if we inferred an invariant that the total supply of fUSDC should not surpass $160$M, the profitability of each round of the exploit would be significantly reduced, making it unable to cover the cost of manipulating the market of Curve. Importantly, both invariants do not affect any normal user’s transaction in histories, making them practical for real-world deployment.
Patch Smart Contracts Post-Deployment: Despite the immutable nature of deployed smart contracts, developers still have different methods to alter their behavior post-deployment, allowing for the addition or modification of invariant guards to shield against future exploits: (1) Upgradable or Modular Contract Design: Upgradable contract standards such as ERC897 (ERC, 2024b) and ERC1167 (ERC, 2024a) incorporate a proxy and an execution contract. The proxy contract delegates function calls to the execution contract, whose address can be modified within the proxy, allowing developers to update their smart contracts after deployment. Similarly, developers can segment a protocol into multiple contracts as different modules. A primary contract interfaces with users, subsequently interact with other modular contracts that handle distinct functionalities including invariant checking. The addresses of modular contracts could be updated in the primary contract. (2) Application Interface Adjustments: For deployed protocols with neither upgradable nor modular contract designs, addressing vulnerabilities or enforcing invariants can still be achieved by launching a revised protocol version and redirecting users through website or application interfaces.
Research Questions: Inspired by these observations and their implications for enhancing smart contract security, we are motivated to explore the following research questions:
RQ1: Given the fact that exploit transactions often exhibit abnormal behaviors, what kinds of invariant guards are most effective at stopping exploit transactions?
RQ2: If an exploit transaction or benign transaction violates invariant guards, how difficult is it for an attacker or a regular user to bypass them?
RQ3: As multiple dimensions of abnormality may be associated with an exploit transaction, how effective is the combination of different invariant guards in preventing exploits?
RQ4: Invariant guards require additional gas at runtime. What are the gas overheads of different types of invariant guards?
RQ5: In terms of enhancing smart contract security, how does this work compare to other state-of-the-art works in smart contract invariant generation or transaction anomaly detection?
4. Invariants
Scope. Our research focuses on the invariants that can be used to distinguish between benign and malicious transactions. Particularly, we focus on the invariants that are broadly applicable to most common DeFi protocols. Invariants that are not for security purposes or highly specific to a single protocol are outside the scope of our study.
Methodology. In our effort to build a comprehensive list of smart contract transaction guards, we carried out an analysis of existing research papers on smart contract security (Liu et al., 2022b; Ghaleb et al., 2023; Rodler et al., 2018; Choi et al., 2021; Breidenbach et al., 2021). Additionally, we conducted a qualitative study on both audit reports and source code of the $63$ audited projects from ConsenSys, a leading smart contract auditing firm, from May 2020 to March 2023 (con, 2023). One author extracted $2181$ enforced invariants from the audit reports and smart contracts under audit by searching for keywords such as “require” and “assert”, after eliminating any duplicates. The author then manually reviewed these invariants to extract templates for pattern matching against remaining uncategorized invariants. This iterative process continued until no new templates could be extracted, and all remaining invariants were also deemed uncategorized for specific reasons. This task took three weeks. Following this, another author reviewed and validated both the categorized and uncategorized invariants for accuracy. In cases of disagreement, a third author was consulted to resolve the issue. This review process lasted two weeks. All three authors have over two years of smart contract security research experience.
Table 1. Qualitative Study Statistics Overview (The table’s left section presents key statistics from the qualitative study. The middle section presents categorized instances across invariant categories. The right section presents instances for various reasons why these invariants remain uncategorized.)
Statistics Count Category Count Reason Count
# Audits 63 Access Control 283 Protocol Specific 1098
# Code Repositories 49 Time Lock 158 Array Length Check 200
# Invariants in Total 2181 Gas Control 2 Byte Operation 44
# Invariants Categoried 826 Re-entrancy 12 Safe Math 13
# Invariants Uncategorized 1355 Oracle Slippage 15
Special Storage 24
MoneyFlow 151
DataFlow 181
The above process resulted in $826$ invariants under $8$ categories, which comprise $37.87\%$ of the overall $2181$ invariants, as shown in Table 1. Access Control and DataFlow are the top and second common categories, respectively. The remaining invariants ($62.13\%$) can not be categorized for various reasons. The most common reason is that invariants are specific to a particular protocol. For example, the invariant “require(validUniswapPath(bAsset))” checks whether “bAsset” is a valid Uniswap path. However, this invariant can only apply to protocols involving Uniswap, thus limiting its applicability. Other uncategorized invariants are used for checking array lengths, byte operations, and arithmetic safety, which target specific data structures or operations. Such invariants of low-level operations are hard to apply as security guards because they are unable to capture the high-level user intentions.
Invariant Templates. Table 2 summarizes the results of our study. In the table, the Category column groups invariant templates based on their application domains, such as Access Control, Time Lock, etc. The ID column assigns a unique identifier to each invariant, while the Name column provides a human-readable description. The Template column contains formal representation of the invariant templates. Specifically, we use $\mathsf{x}$ to denote a contract state record maintained by invariant templates. We use $\mathsf{r}$ to represent a local variable and _?_?\_?_ ? as the undetermined parameter to be inferred. The Parameter column shows the type of the undetermined parameter. The References column lists the academic or industry sources of each invariant template.
Table 2. Invariants. (We use $\mathsf{x}$ to denote a contract state variable, $\mathsf{r}$ to denote a local variable, and _?_?\_?_ ? to denote a hole in the template to fill during the synthesis. |_|_|\_|| _ | denotes the absolute value.)
Category ID Name Template Parameter References
Access Control EOA onlyEOA $\mathsf{msg.sender}=\mathsf{tx.origin}$ (Liu et al., 2022b; Ghaleb et al., 2023)
SO isSenderOwner $\mathsf{msg.sender}=\mathsf{owner?}$ address
SM isSenderManager $\mathsf{msg.sender}=\bigcup_{i=1}^{n?}\ \mathsf{mgr_{i}?}$ addresses
OO isOriginOwner $\mathsf{tx.origin}=\mathsf{owner?}$ address
OM isOriginManager $\mathsf{tx.origin}=\bigcup_{i=1}^{n?}\ \mathsf{mgr_{i}?}$ addresses
Time Lock SB isSameSenderBlock $\mathsf{x_{entrySdrBlk}}\neq\mathsf{r_{exitSdrBlk}}$ (idl, 2023)
OB isSameOriginBlock $\mathsf{x_{entryOrgBlk}}\neq\mathsf{r_{exitOrgBlk}}$
LU lastUpdate $\mathsf{r_{curtBlk}}-\mathsf{x_{lstBlk}}\geq\mathsf{nbBlks?}$ Integer (mst, 2023; fei, 2023b, a)
Gas Control GS GasStartUpperBound $\mathsf{gasStart}\leq\mathsf{gas?}$ Integer motivated by Section 3
GC GasConsumedUpperBound $\mathsf{gasStart}-\mathsf{gasEnd}\leq\mathsf{gas?}$ Integer
Re-entrancy RE nonReEntrant $\mathsf{x_{lock}}=\mathsf{true}$ (Rodler et al., 2018)
Oracle Slippage OR OracleRange $\mathsf{pr_{LB}?}\leq\mathsf{r_{newPr}}\leq\mathsf{pr_{UB}?}$ Integer (dfo, 2023b)
OD OracleDeviation $|(\mathsf{r_{newPr}}-\mathsf{x_{oldPr}})/\mathsf{x_{oldPr}}|\leq\mathsf{pr_{Dev}?}$ Integer (dfo, 2023b; Breidenbach et al., 2021)
Special Storage TSU TotalSupplyUpperBound $\mathsf{x_{totSup}}\leq\mathsf{totSup?}$ Integer (Aav, 2023; dfo, 2023a)
TBU TotalBorrowUpperBound $\mathsf{x_{totBor}}\leq\mathsf{totBor?}$ Integer (Aav, 2023; dfo, 2023a)
MoneyFlow TIU TokenInUpperBound $\mathsf{r_{tokenIn}}\leq\mathsf{v?}$ Integer (bal, 2023; dfo, 2023a)
TIRU TokenInRatioUpperBound $\mathsf{r_{tokenIn}}\leq\mathsf{v?}$ Integer (bal, 2023)
TOU TokenOutUpperBound $\mathsf{r_{tokenOut}}/\mathsf{b_{token,adr}}\leq\mathsf{v?}$ Integer (bal, 2023; dfo, 2023a)
TORU TokenOutRatioUpperBound $\mathsf{r_{tokenOut}}/\mathsf{b_{token,adr}}\leq\mathsf{v?}$ Integer (bal, 2023)
DataFlow MU MappingUpperBound $\mathsf{map?[\mathsf{index?}]}\leq\mathsf{v?}$ Integer (Choi et al., 2021)
CVU CallValueUpperBound $\mathsf{msg.value}\leq\mathsf{v?}$ Integer
DFU DataFlowUpperBound $\mathsf{var?}\leq\mathsf{v?}$ Integer
DFL DataFlowLowerBound $\mathsf{var?}\geq\mathsf{v?}$ Integer
4.1. Access Control
Many research papers have conducted extensive studies on the access control (Liu et al., 2022b; Ghaleb et al., 2023). Access control governs the privileges associated with the transaction’s sender and origin, dictating which addresses are authorized to invoke specific smart contract functions.
Note that transaction’s sender and origin could be different in Ethereum. The sender is the address which invokes the contract function, while the origin is the address who initiates the entire transaction. For example, if user address $a$ calls contract $b$ which in turn calls contract $c$, during the execution of $c$, the sender address is $b$ while the origin address is $a$.
onlyEOA (EOA): This template verifies that the transaction’s origin matches the sender’s address, thereby confirming it was initiated from an externaly owned user address (i.e., EOA address) rather than a contract address. The intuition of this invariant template is that many attack strategies involves multiple sophisticated interactions and therefore attackers often have to write their own contracts. This template can neutralize such attack strategies.
isSenderOwner (SO) and isOriginOwner (OO): These templates restrict function execution to a predefined address ($\mathsf{owner?}$) that are registered as owners.
isSenderManager (SM) and isOriginManager (OM): These templates only allow function calls from a set of predefined manager addresses $\mathsf{mgr_{i}?}$.
The access control invariants are typically inserted at the beginning of non-read-only functions to immediately halt unauthorized attempts to alter contract state.
4.2. Time Lock
The Time Lock category of invariants serves as a temporal gating mechanism for smart contract functions. This category contains three invariants.
isSameSenderBlock (SB) and isSameOriginBlock (OB): These templates limit the ability to execute specific paired functions within the same block by the same sender or origin. For example, to inhibit the same sender or origin address from invoking both the deposit and withdraw functions consecutively within a single block. The intuition is that normal users are unlikely to initiate multiple interactions with the same function in a few seconds, while malicious attackers often use iterative loops to drain funds from a victim contract. To implement these invariants, a state variable $\mathsf{x_{entrySdrBlk}}$ (resp., $\mathsf{x_{entryOrgBlk}}$) stores a hashed combination of the transaction sender (resp., origin) address and the current block number upon entry into a function (e.g., deposit). In the exit function (e.g., withdraw), this stored value is compared against a freshly computed hash, stored in $\mathsf{r_{exitSdrBlk}}$ (resp., $\mathsf{r_{exitOrgBlk}}$), to ensure that they differ. These two invariants are designed to be updated at the entry point of enter functions, i.e., functions that accept tokens from users. Then verified at the start of exit functions, i.e., functions that are responsible for disbursing tokens back to users.
lastUpdate (LU): These template moderates the frequency with which a given function can be invoked. It inserts guard at the beginning of non-read-only functions to mandate that a specified number of blocks, denoted as $\mathsf{nbBlks?}$, must elapse between two consecutive calls to the same function. To enforce this, the state variable $\mathsf{x_{lstBlk}}$ captures the timestamp of the last block where the function was invoked. Subsequent calls to the function check this stored timestamp against the current block timestamp. The difference must meet or exceed the $\mathsf{nbBlks?}$ threshold.
4.3. Re-entrancy
The Re-Entrancy class of invariants tackles re-entrancy vulnerabilities in smart contracts. Represented by a single invariant template, nonReEntrant (RE), this category utilizes a state variable $\mathsf{x_{lock}}$ as a lock to prevent a transaction from entering a set of key functions of a contract more than once. $\mathsf{x_{lock}}$ will be set to $True$ when a function is invoked and reset to $False$ when a function returns. The RE guard is usually placed at the beginning of enter and exit functions of a contract to effectively mitigate re-entrancy risks.
4.4. Gas Control
We propose the Gas Control category of invariants, motivated by the Harvest example. The intuition is that malicious attacks tend to have significantly more complicated logic to consume a large amount of gas. This class consists of two invariants: GasStartUpperBound (GS) and GasConsumedUpperBound (GC). As illustrated in Table 2, the GS invariant sets an upper limit on the remaining gas at the entry point of a function, using the variable $\mathsf{gasStart}$, whereas the GC invariant sets an upper bound on the total gas consumed within the function by comparing the remaining gas at the entry and exit point of a function, $\mathsf{gasStart}$ and $\mathsf{gasEnd}$. These invariants are designed to be placed at the beginning and end of non-read-only functions.
4.5. Oracle Slippage
The Oracle Slippage category mitigates risks tied to price oracles in decentralized finance (DeFi) applications. The intuition of templates in this category is to detect potential price manipulation by malicious attacks. This class includes two invariant templates: OracleRange (OR) and OracleDeviation (OD). The OR template enforces a bounded range for the oracle prices. It utilizes two parameters, $\mathsf{pr_{LB}?}$ and $\mathsf{pr_{UB}?}$. The OD template enforces a specific percentage deviation limit between the current and last price provided by the oracle. The parameter $\mathsf{pr_{Dev}?}$ is employed to define a permissible deviation rate. These invariants are usually inserted right after the oracle is called.
4.6. Special Storage
The Special Storage class of invariant templates is concerned with constraining global storage variables that are crucial to the contract’s state or logic. The intuition of this is after an exploit, the contract’s state variables are often in an abnormal state. Thus, by constraining the state variables, we can prevent the exploit. This class features two main invariants: TotalSupplyUpperBound (TSU) and TotalBorrowUpperBound (TBU). The TSU invariant imposes an upper bound (denoted as $\mathsf{totSup?}$) on the contract’s totalSupply variable, while TBU sets a ceiling (denoted as $\mathsf{totBor?}$) for the contract’s totalBorrow variable which represents the total amount that can be borrowed from the contract. To preserve the integrity of these important state variables, these invariants are inserted at the functions that could modify the total supply or total borrow balances.
4.7. Money Flow(also called Token Flow)
The MoneyFlow class focuses on the flow of tokens within the smart contract, particularly for functions involving token deposits and withdrawals. The intuition of these templates is that malicious transactions tend to cause abnormally large amount of digital asset movement.
TokenInUpperBound (TIU) and TokenOutUpperBound (TOU): This template caps the number of tokens flowing into or out of the contract each time by using an integer parameter $\mathsf{v?}$.
TokenInRatioUpperBound (TIRU) and TokenOutRatioUpperBound (TORU): These templates constrain the ratio of tokens flowing into or out of the contract, in relation to the contract’s current token balance. They also employ an integer parameter $\mathsf{v?}$.
To ensure effective governance of money flow, these invariants are placed within functions that handle token transfers. The TIU and TIRU invariants are applied right before a token deposit, while the TOU and TORU invariants are applied before a token withdrawal.
4.8. DataFlow
Smartian (Choi et al., 2021) leverages dynamic taint analysis to detect whether a block state can affect an ether transfer. We extend their work to include all data flows affecting both ether and ERC20 token transfers. This allows us to set constraints on values that could potentially be controlled by an attacker to manipulate transfer amounts. The DataFlow category is subdivided into four specific invariants, each designed to address a particular type of variables in the data flow.
MappingUpperBound (MU): This invariant focuses on values stored in the contract’s mapping data structure, often representing a user’s property(e.g., a user’s share in a liquidity pool). To constrain such user-specific values, we introduce a parameter $\mathsf{v?}$ to set an upper limit on these mappings.
CallValueUpperBound (CVU): Call values signify the amount of ether transferred during a function call. Because these values directly affect the contract’s ether balance, we list it as a separate invariant. We employ a parameter, denoted as $\mathsf{v?}$, to cap the incoming ether to mitigate risks of abnormal or malicious deposits.
DataFlowUpperBound (DFU) and DataFlowLowerBound (DFL): These invariants apply to all other data flow variables, whether derived from external calls, storage loads, or calldata. To regulate these variables, we use a parameter $\mathsf{v?}$, setting either upper or lower bounds on these values to thwart unauthorized manipulations.
The above invariants are inserted at the locations where data flow variables are first read. This is often in functions that initiate token transfers. These invariants helps the contract ensure that every value used for the calculation of token transfers is within the normal range.
5. Trace2Inv
We present Trace2Inv, our proposed framework to infer concrete invariants as introduced in Section 4 for a given contract by analyzing its historical transaction traces. As shown in Figure 2, Trace2Inv is structured into three modules: trace parser, invariant-related data extraction, and invariant generation. The second module consists of three sub-modules: invocation tree analysis, type inference, and dynamic taint analysis.
Figure 2. An Overview of Trace2Inv
5.1. Trace Parser
A transaction’s trace data, denoted as structLogs, contains a sequence of executed EVM instructions and each instruction is a six-item tuple $\mathsf{\langle pc,op,gasLeft,gasCost,stack,memory\rangle}$ that includes the current program counter, the EVM opcode to execute, the amount of the remaining gas and the gas consumed by current opcode execution, the full view of current EVM stack and memory. In trace parser, we reconstruct the functional context information from structLogs as an invocation tree where each node represents an external function call and the node hierarchy reflects the call-chain relationship. In particular, each tree node is a seven-item tuple $\mathsf{\langle addr,func,args,ret\_data,ins,gasEntry,gasExit\rangle}$ that records the current contract address, function name, corresponding arguments, returned data, the set of executed EVM instructions belonging to the function, and the amount of the remaining gas at the entry and exit points of the function. The invocation tree also contains metadata of the transaction, such as the transaction hash, the block number, and the origin address. The trace parser leverages the target contract address to isolate a selected trace segment corresponding to the target contract’s execution. This invocation tree and the segmented trace data are then passed to the next module for further analysis.
5.2. Invocation Tree Analysis
Invocation tree analysis extracts invariant-related data exclusively from the invocation tree, sufficient to collect data for access control, time lock, gas control, re-entrancy, oracle, and money flow invariants. For instance, for the sender opcode, its results can be extracted from the invocation tree by searching for parent node of the node of target contract. Moreover, the invocation tree can also be used to identify locations of re-entrancy by capturing the nested and recursive calls to the target contract. Oracle values are also obtained by traversing the invocation tree to locate calls to the oracle. For money flow invariants, the required values can also be directly read from function calls of transfering Ether/ERC20.
5.3. Type Inference
Type inference in Trace2Inv infers the type of storage slot when it is accessed. It is essential for extracting data relevant to both special storage and data flow invariants, as accurately decoding storage accesses with the correct type is necessary for these invariant categories. Decoding storage slots is straightforward when they are listed in the contract’s storage layout, which is typically the case for special storage invariants. However, the challenge arises with complex data structures like mapping, often used in data flow invariants. These structures may reference storage slots not explicitly present in the contract’s storage layout, which instead are computed through operations such as sha3 and arithmetic functions. To solve this issue, Trace2Inv maintain a preimage dictionary to track the key’s computation. When Trace2Inv encounters a sha3 opcode, the mapping positions and slots are recorded. In Solidity, the first $32$-bytes represent the mapping slot, and the next $32$-bytes serve as the mapping position. In Vyper, these roles are reversed. Anytime a 64-byte sha3 hash is encountered, both its hash value and origin are recorded for future reference. When an sload operation has a key that is not present in the storage layout, the preimage dictionary is consulted. We recursively trace back the computation steps until we identify a mapping data structure that is in the storage layout. Utilizing the types of the mapping data structure, we can infer the type of storage slot accessed.
5.4. EVM-level Dynamic Taint Analysis
Dynamic taint analysis in Trace2Inv operates at the EVM opcode level to track data sources. It is used to extract data for data flow invariants. Given a target contract, a corresponding trace segment, and the invocation tree of the transaction, the taint analyzer collects all accessible information pertaining to that contract. The analyzer also infers the data types of recorded tainted data or taint sources. It accomplishes this by utilizing storage layouts and function ABIs thereby providing a comprehensive, type-aware taint analysis tailored for smart contracts.
Table 3. Instructions Defined as Sources and Sinks.
Category Opcodes and Locations
Sources External Address Variables balance, extcodesize, extcodecopy, extcodehash,
Execution Context Variables origin, caller, address, codesize, selfbalance, pc, msize, gas
Call Data Variables callvalue, calldataload, calldatasize, calldatacopy
Return Data Variables returndatasize, returndatacopy
Block Variables blockhash, coinbase, timestamp, number, prevrandao, gasprice, gaslimit, chainid,
Storage sload(untainted)
Sinks ether transfer address.call{value: uint256}
ether transferFrom callvalue
ERC20 transfer ERC20.transfer(address,uint256), ERC20 = token
ERC20 transferFrom ERC20.transferFrom(address1,address2,uint256), ERC20 = token, address2 = this
Taint Sources and Sinks. Table 3 lists the EVM instructions that our dynamic taint analyzer identifies as sources and sinks for taint propagation. The table is divided into two categories: Sources and Sinks. In the Sources category, we outline various sub-categories of taint sources, which include external address, execution context, call and return data, block variables, and storage. Opcodes like balance, extcodesize, and sload are some of EVM opcodes that load new taint sources into the stack or memory. They are key to the taint propagation as they introduce data that could potentially influence other data points or outcomes in the contract.
[…]
Bit Level Taint Propagation. Our taint analyzer maintains three distinct taint trackers: the stack tracker, the memory tracker, and the storage tracker. Initially, all these trackers are empty. Each tracker records taint information at the bit level, enabling granular analysis. For instance, sload and sstore opcodes can only read and write $32$-byte chunks, but the taint status is stored for each individual bit. Our propagation of taints follows a set of rules based on prior work (Ghaleb et al., 2022):
R1: A value derived from one or more operands becomes tainted if any of the operands is tainted.
R2: For sload, if the storage slot loaded is tainted, the $32$-byte stack entry receives the corresponding taint information. If not, the sload acts as a new taint source and taints the stack entry. For sstore, the $32$-byte entry in the storage tracker is overwritten with the taint information from the stack.
R3: For instructions that read data from memory (e.g., mload), the result value is tainted if the read data is tainted. The same logic applies for data loaded from storage using the sload opcode.
R4: In external calls, if the arguments in memory are tainted, the return data will also be tainted.
5.5. Invariant Generation
Using the extracted data by the previous module, the invariants’ inference module uses the templates provided in Table 2 to generate concrete invariants where all holes are filled with concrete values. The synthesis heuristic of generating invariants varies based on their category:
Hypothesis Testing: For EOA, SB, OB, and RE invariants, which act as assumptions regarding the contract’s behaviors, no parameters need to be learned. If no data points violate these invariants (i.e, all transactions in the training set satisfy the invariant), they are then directly applied to the contract. Otherwise, the violated invariant is not applied.
Role-based Set Inference: For address-based invariants such as SO, SM, OO, and OM, we adopt a set-based heuristic. We analyze the set of senders or origins associated with each function call. If the size of this set exceeds a certain threshold (greater than $1$ for owners or $5$ for managers), we consider it as a violation, and the corresponding invariant is not applied to that function.
Bound Deduction: Invariants in the categories of gas control, oracle slippage, special storage, and money flow require learning the bounds of certain integer parameters. For those, the maximum and minimum values are chosen among the collected data points, provided the data contains at least two distinct values. Particularly for the oracle slippage invariant (OR), an additional tolerance of $20$% is included for the upper and lower bounds, in accordance with prior research (Tolmach et al., 2021). For TIRU and TORU invariants, we filter outliers using a z-score threshold of $3$.
Hybrid: Lastly, for the LU invariant, which deals with the block gap between calls to the same function, we calculate the smallest block gap among the data points. If any two data points have a block gap of zero, the invariant is not applied to the corresponding function. Otherwise, the smallest block gap is used as the parameter for the invariant.
5.6. Trace2Inv Implementation
Trace2Inv’s input, transaction trace data, can be obtained from Ethereum via its API debug_traceTransaction. Within the Trace Parser module, Trace2Inv queries EtherScan (Eth, 2020) to gather additional transaction metadata not directly available in the trace data, such as block number and transaction origin. This metadata provides the execution context crucial for generating certain invariants, like EOA and LU. Additionally, Trace2Inv also queries EtherScan (Eth, 2020) to fetch the contract’s source code. The code is then compiled locally using the appropriate versions of the Solidity (sol, 2018) or Vyper (vyp, 2018) compilers to obtain the contract’s function ABI. With this ABI, Trace2Inv leverages Slither (sli, 2021) to decode function names, their arguments, and return values present in the trace. Within Invariant Related Data Extraction Module, Trace2Inv compiles source code of target smart contract, and reads the compiler’s output to obtain its storage layout. However, some old versions of Solidity and Vyper compilers do not support this functionality. In such cases, Trace2Inv fetches the storage layout from EVM Storage (EVM, 2024).
Several optimizations are incorporated into Trace2Inv implementation to enhance its performance. Firstly, when fetching trace data from an Ethereum archive node, we optimize the process by using batching RPC requests. This significantly reduces the overhead associated with individual API calls. Secondly, we employ parallelization techniques to speed up the process of fetching and parsing trace data. Specifically, multi-processing is used to concurrently handle different segments of trace data, converting them into summaries in a more time-efficient manner. Thirdly, caching is utilized to further optimize performance; the system caches query results from EtherScan, archive nodes, and compilers. This minimizes redundant executions and API calls, thereby accelerating the overall analysis process.
6. Evaluation
Table 4. Benchmarks. (Exploit: the first exploit transaction during the incident. FL: whether the exploit transaction uses flash loan. Contracts: the victim contracts involved in the exploit. Type: the type of the contract, either interface (I) or payload (P). PL: the programming language of the contract, either Solidity (S) or Vyper (V). History: the number of transactions in the history of the contract up to the exploit transaction.)
    Kind Victim Protocol Root Cause Date Loss Exploit FL Contracts Type PL History
    Bridges RoninNetwork keys compromised 22/03/29 624M (Etherscan, 2023u) RoninNetwork I+P S 95345
HarmonyBridge keys compromised 22/06/24 100M (Etherscan, 2023h) HarmonyBridge P S 32149
Nomad
zero hash as
a valid root
22/08/01 152M (Etherscan, 2023m) Nomad I+P S 15630
PolyNetwork hash collision 21/08/10 611M (Etherscan, 2023p) PolyNetwork I+P S 44509
    Lending bZx2 oracle manipulation 20/02/18 630K (Etherscan, 2023b) bZx2 I+P S 711
Warp oracle manipulation 20/12/17 8M Warp P S 148
(Etherscan, 2023y) Warp_I I S 31
CheeseBank oracle manipulation 20/11/06 3.3M CheeseBank_1 I+P S 615
CheeseBank_2 I+P S 593
(Etherscan, 2023c) CheeseBank_3 I+P S 557
InverseFi oracle manipulation 22/06/16 1.26M (Etherscan, 2023l) InverseFi I+P S 7590
CreamFi1 cross contract re-entrancy 21/08/30 18M CreamFi1_1 I+P S 5
(Etherscan, 2023d) CreamFi1_2 I+P S 58184
CreamFi2 oracle manipulation 21/10/27 130M CreamFi2_1 I+P S 1270
CreamFi2_2 I+P S 898
CreamFi2_3 I+P S 261
(Etherscan, 2023e) CreamFi2_4 I+P S 98
RariCapital1 read-only re-entrancy 21/05/09 10M (Etherscan, 2023r) RariCapital1 I+P S 667
RariCapital2 cross contract re-entrancy 22/04/30 80M RariCapital2_1 I+P S 614
RariCapital2_2 I+P S 752
RariCapital2_3 I+P S 404
(Etherscan, 2023s) RariCapital2_4 I+P S 776
XCarnival logic error 22/06/26 3.87M (Etherscan, 2023z) XCarnival I+P S 342
    Yield-Earning Harvest1 oracle manipulation 20/10/26 33.8M (Etherscan, 2023i) Harvest1 I+P S 2050
Harvest2 oracle manipulation 20/10/26 (Etherscan, 2023j) Harvest2 I+P S 2161
ValueDeFi oracle manipulation 20/11/14 7.4M (Etherscan, 2023w) ValueDeFi I+P S 295
Yearn forced investment 21/02/04 11M Yearn P S 672
(Etherscan, 2024) Yearn_I I S 26688
VisorFi re-entrancy 21/12/21 8.2M (Etherscan, 2023x) VisorFi I+P S 1693
UmbrellaNetwork underflow bug 22/03/20 700K (Etherscan, 2023v) UmbrellaNetwork I+P S 59
PickleFi access control 20/11/21 20M (Etherscan, 2023o) PickleFi I+P S 5439
    Others Eminence logic error 20/09/29 7M (Etherscan, 2023g) Eminence I+P S 20589
Opyn logic error 20/08/04 371k (Etherscan, 2023n) Opyn I+P S 67
IndexFi logic error 21/10/15 16M (Etherscan, 2023k) IndexFi I+P S 20641
RevestFi re-entrancy 22/03/27 11.2M RevestFi P S 1635
(Etherscan, 2023t) RevestFi_I I S 1463
DODO access control 21/03/08 700K (Etherscan, 2023f) DODO I+P S 42
Punk access control 21/08/10 8.9M Punk_1 I+P S 28
Punk_2 I+P S 42
(Etherscan, 2023q) Punk_3 I+P S 37
BeanstalkFarms flashloan assisted commit 22/04/16 182M BeanstalkFarms P V 5785
(Etherscan, 2023a) BeanstalkFarms_I I S 306
In this section, we aim to empirically answer the research questions raised in Section 3 by applying invariant guards to real-world exploits on Ethereum Blockchain. We systematically collected economic exploit incidents that cost greater than 300K USD financial loss from February 14, 2020, to August 1, 2022 on Ethereum Blockchain. Our benchmark is compiled from a diverse set of sources including academic publications (Chen et al., 2022; Qin et al., 2021), industry databases (SlowMist, 2023; BlockSec, 2023), and open-source GitHub repositories (Contributors, 2023a, b). It is worth noting that we exclude from our benchmarks any hacks targeting individual user wallets, as these are primarily the result of private key leakage, rather than protocol vulnerabilities. We also exclude hacks where the victim contracts are close-source, as our manual analysis requires the source code of the victim contracts. Note that Trace2Inv can also be applied to close-source contracts, as long as their function ABIs and storage layout are available.
Table 4 presents the benchmark dataset in our study. It comprises 27 hacks which cumulatively resulted in financial losses exceeding 2 billion USD. The column FL denotes whether the exploit involves flash loans, which require atomicity for the hack transactions. The column Type denotes the type of victim contracts. For each exploit, two types of victim contracts are manually identified: payload (P) refers to the protocol contract that eventually transfers abnormal amounts of tokens out of the protocol, and interface (I) refers to the contract that is directly invoked by users to initiate this abnormal transfer. The History column signifies the length of the transaction history up to the exploit transaction for each victim contract.
6.1. RQ1: Effectiveness of Smart Contract Invariants
Experiment. In our first experiment, we utilize $23$ pre-defined invariant templates, as detailed in Section 4, to dynamically infer invariants from the transaction histories of $42$ different victim contracts, using the invariant generation methods as described in Section 5.5. We divide each contract’s transaction history into two distinct sets: 70% of the transactions are allocated for training set, while the remaining 30% are used as the test set.
To validate the effectiveness of the invariants dynamically inferred, we employ the transaction trace data in the test set for evaluation. Utilizing the same parser and dynamic taint analyzer, we obtain the invocation tree and data points pertinent to the particular invariant for each transaction. With these, we can evaluate whether a transaction violates any of the invariant guards in place. If a transaction is blocked by these invariant guards, it serves as a positive example for the effectiveness of the invariants. The validation process discriminates between different kinds of positives. Specifically, if the exploit transaction is successfully blocked by the invariant, it is categorized as a True Positive. On the other hand, any non-exploit transactions blocked are counted as False Positives.
Table 5. Summarized Results of Invariants Effectiveness Evaluation.
AccessControl TimeLock GasCtrl Re Oracle Storage MoneyFlow DataFlow
    EOA     SO     SM     OO     OM     SB     OB     LU     GS     GC     RE     OR     OD     TSU     TBU     TIU     TOU     TIRU     TORU     MU     CVU     DFU     DFL
# Contracts Applied(42) 42 39 22 39 25 33 33 37 41 41 40 11 11 21 16 34 34 28 28 11 7 33 33
# Contracts Protected(42) 23 6 7 5 9 11 13 12 30 23 2 7 6 8 12 9 22 5 22 1 2 23 1
# Hacks Blocked(27) 15 4 6 3 8 9 11 10 18 15 2 5 4 7 6 8 17 5 15 1 2 18 1
Average FP(%) 0.2 0.6 1.8 0.5 2.6 0 0 2.4 2.4 2.6 0 22.3 4.6 7.9 12.6 0.7 0.4 0.8 1.4 1.5 0.7 1.6 0.9
Results. Table 5 provides a comprehensive summary of the effectiveness evaluation for the invariants applied across various contract categories. The table lists several key metrics: the number of contracts to which each type of invariant is applied, the number of contracts successfully protected by these invariants, the number of hacks blocked, and the Average False Positive rate. Among these, the row # Hacks Blocked stands out as the most crucial metric as it directly measures the capability of each invariant to block financial exploits. The Average FP rate is also an important metric as it quantifies the potential impact on regular users, reflecting the trade-off between security and usability.
The applicability of the invariants varies across different categories. Access Control, Time Lock, Gas Control, MoneyFlow, and DataFlow are universally applicable, protecting a broad range of contracts and blocking numerous hacks. On the contrary, categories such as ReEntrancy, Oracle, and Storage have narrower scopes, applicable only to specific types of contracts. For instance, the ReEntrancy invariant we studied is effective only against common single-contract reentrancy attacks. Other attack types, such as read-only or cross-contract reentrancy seen in CreamFi1, RariCapital1, and RariCapital2, require more specialized invariants and are left as future work.
For true positives, in each category of Access Control, Time Lock, Gas Control, Data Flow, and Money Flow, there is a standout invariant that proves most effective at blocking hacks: EOA for Access Control, OB for Time Lock, GS for Gas Control, TOU for Money Flow, and DFU for Data Flow. These invariants block the highest number of hacks in their respective categories.
For false positives, EOA, OB and TOU have an average FP rate below 0.4%, while GS and DFU also have a low FP rate below 2.4%. This low rate indicates that these invariants have a small impact on regular user transactions, thereby making them practical for real-world deployment. The elevated false positive rates observed for OR, TSU, and TBU are primarily because of the fluctuating nature of oracle values, total supply, and total borrow. Using upper-bound or range-based invariants for these categories could inadvertently block all transactions once these values exceed a certain threshold.
Answer to RQ1: EOA in Access Control, OB in Time Lock, GS in Gas Control, TOU in Money Flow, and DFU in Data Flow shows the highest efficacy in blocking hacks while having a low false positive rate.
6.2. RQ2: Study of False Positives and True Positives
Case Studies. In our second research question (RQ2), we explore the bypassability of the invariants for both malicious hackers and normal users. Hackers will be informed when the invariants are deployed in the target contract from the source code or the bytecode. It is natural to ask whether malicious hackers can bypass the invariants and still gain profit if they realize the existence of such invariant guards. We manually analyze every exploit transaction blocked by each invariant (true positives) to check its bypassability. For each case, we assign one of three categories: C1: the exploit is entirely blocked, and the hacker can no longer gain any profit; C2: the exploit is partially blocked, resulting in significantly reduced profits for the hacker; and C3: the hacker can still achieve profits similar to historical data, with some adjustments to their exploit code.
We also manually analyze the false positives generated by our invariants to assess their impact on regular, benign users. For this, we sample up to 10 transactions from the false positives for each invariant and manually evaluate their bypassability. We operate under the assumption that regular users have the option to split transactions with large parameters into transactions with smaller parameters (via DApp’s front-end), lower the gas for their transactions (via a wallet software), or simply wait for some time to transact again after their transaction previously blocked. Based on these criteria, we categorize the false positives into three groups: D1: the transaction is completely blocked and cannot be bypassed through simple means; D2: the transaction can be bypassed by breaking it down into smaller transactions; and D3: users can bypass the transaction by reducing the gas or waiting for some time. For both true positives and false positives, two authors independently labeled the bypassability results with the third one to resolve the divergence of views.
Table 6. Bypassability Results of Hacks Blocked (TPs) and Sampled Normal Transactions Blocked (FPs).
Access Control Time Lock GasCtrl Re Oracle Storage MoneyFlow DataFlow
    EOA     SO     SM     OO     OM     SB     OB     LU     GS     GC     RE     OR     OD     TSU     TBU     TIU     TOU     TIRU     TORU     MU     CVU     DFU     DFL
    TPs C1 15 4 6 3 8 2 10 5 0 15 2 3 2 2 3 7 12 2 7 1 1 12 1
C2 0 0 0 0 0 0 0 0 1 0 0 2 2 5 3 0 0 2 8 0 0 0 0
C3 0 0 0 0 0 7 1 5 17 0 0 0 0 0 0 1 5 1 0 1 1 6 0
    FPs D1 10 10 10 10 10 0 0 1 0 10 0 10 10 10 10 0 0 0 1 10 0 4 10
D2 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 10 10 10 9 0 8 6 0
D3 0 0 0 0 0 0 0 9 10 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Results. Table 6 offers a comprehensive view of how both attackers and normal users can potentially bypass the invariant guards. The table is divided into two major rows: True Positives (TPs) and False Positives (FPs). For TPs, we have further categorized the effectiveness of the invariant guards as C1, C2, and C3, signifying the level of success the hacker has in bypassing the guard. For FPs, we use tags D1, D2, and D3 to illustrate how easily normal users can circumvent these guards.
Some invariant categories can easily be bypassed by both hackers and normal users. For example, GS, though it blocks the most hacks in RQ1, can be easily bypassed by changing the gas passed to the specific function call of the target contract. Some other invariants, such as GC, cannot be bypassed by either hackers or normal users. EOA also fall into this category. OB interestingly has no false positives, but in practice it is very hard to be bypassed by normal users.
For certain invariants such as TOU and DFU, there is a dichotomy where they block exploit transactions effectively but normal users can still bypass them. They make certain exploits impossible by preventing hackers from reaching specific contract states required for profitability. Regular users, who usually do not attempt to manipulate contract states for exploitative gains, can often bypass these invariants by simply splitting their larger transactions into smaller ones. Though this results in more transactions and higher gas costs, it does not impede their primary objectives.
Answer to RQ2: Most of invariants behave similarly for both hackers and normal users, either being easily bypassed (such as GS) or not bypassable at all (such as EOA, OB, GC). Some invariants could block hackers while allowing normal users to circumvent them (such as TOU and DFU).
6.3. RQ3: Effectiveness of Combination of Invariants
Experiment. In RQ3, we look into the effectiveness of various invariant combinations in safeguarding smart contracts. Building on insights from RQ1 and RQ2, we identify a set of five effective invariants (EOA, SO, GC, TOU, and DFU), replacing GS with GC considering that it can be easily bypassed. Our aim is to explore whether these invariants, each effective in its own domain, can be combined to provide a more robust defense against contract exploits while maintaining a low false positive rate.
After investigating the exploits blocked by each of these five invariants in RQ1, we have observed that (1) EOA and GC can uniquely block 2 and 1 hacks, respectively; (2) no exploits can uniquely be blocked by OB, TOU or DFU; (3) all exploits blocked by MFU can be blocked by DFU.
We then designed an experiment that combines four invariants—EOA, GC, OB, and DFU—each chosen for its efficacy in blocking hacks and resistance to bypass. We consider two logical operators: conjunction ($\land$) and disjunction ($\lor$). By enumerating all logical combinations of invariants up to a length of $4$, we can evaluate their collective True Positives (TPs) and False Positives (FPs) based on two metrics: (1) their ability to block the maximum number of hacks (2) their ability to block the maximum number of hacks while maintaining a false positive rate below 1%.
Results.
Table 7. Validation Results of the Best Combined Invariants.
EOA $\land$ GC $\land$ DFU EOA $\land$ (OB $\lor$ DFU)
# Contracts Applied 42 C1 20 # Contracts Applied 42 C1 18
# Contracts Protected 35 C2 0 # Contracts Protected 31 C2 0
# Hacks Blocked 23 C3 3 # Hacks Blocked 20 C3 2
Average FP rate(%) 3.99 D1 8 Average FP rate(%) 0.32 D1 10
D2 2 D2 0
D3 0 D3 0
Table 7 shows the validation results of the best combined invariants based on metrics (1) and (2), using the same evaluation methodology as described in RQ1 and RQ2.
The combination EOA $\land$ GC $\land$ DFU emerges as the most effective one according to the metric (1). This result is intuitive, as this combined invariant effectively reverts a transaction when any one of EOA, GC, or DFU is violated. Our early observation that EOA and GC can uniquely block 2 and 1 hacks, also justifies their inclusion in this composite invariant. However, this comes at the expense of a higher false positive rate of 3.99%.
The combination EOA $\land$ (OB $\lor$ DFU) maintains an impressively low false positive rate of 0.32%, making it superior based on the metric (2). Its efficacy is due in part to OB’s inherent low false positive rate and to the complementary nature of OB and DFU in the context of token transfer functions. Both combined invariants are better at stopping hacks than individual invariants. This is because they can work together in different parts of the smart contract code, making it more likely they will catch harmful actions. Sometimes a function in the contract may only be protected by one invariant, if other invariants are not applicable to this function. In those cases, both hackers and normal users may still find a way to bypass the security measures.
Answer to RQ3: The invariant guards studied and generated by Trace2Inv are complimentary and their combinations are promising to be more effective on contract protection and hack prevention with lower false positive rate.
6.4. RQ4: Gas Overhead of Invariant Guards
Experiment. In our next research question, RQ4, we examine the gas overhead incurred by the deployment of individual and combined invariants. We select four benchmark contracts that represent different kinds of protocols and programming languages. The target smart contract is instrumented with the generated invariants studied in RQ3 and then compiled by either a Solidity or Vyper compiler. This process is carried out for both the original and the instrumented versions of the contract. The compiler returns an estimate of the gas consumption for each function in the contract, allowing us to calculate the gas overhead for each inserted invariant by comparing the two versions. Subsequently, we replay all transactions within the test set on these instrumented contracts. For each transaction, we add the corresponding gas overhead when the transaction reaches where the invariants are inserted.
The total gas overhead is calculated as, Total Gas After InstrumentationTotal Gas Before InstrumentationTotal Gas Before $\frac{\text{Total Gas After Instrumentation}-\text{Total Gas Before Instrumentation}}{{\text{Total Gas Before Instrumentation}}}.$ This provides a quantitative measure of the computational burden imposed by the invariant guards, aiding in the cost-benefit analysis of their deployment.
Table 8. Runtime Gas Overhead (%) of Different Types of Invariant Guards.
Kind Benchmark Compiler EOA OB GC DFU EOA $\land$ GC $\land$ DFU EOA $\land$ (OB $\lor$ DFU)
Bridge HarmonyBridge Solidity 0.5.17 0.04 0.52 0.02 0.01 0.07 0.56
Lending Harvest1 Solidity 0.5.18 0 1.46 0.01 0.01 0.02 1.46
Yield-Earning CreamFi2_1 Solidity 0.5.17 0 2.55 0 0 0.01 2.55
Others BeanstalkFarms Vyper 0.2.8 0 0.53 / 0 0 0.53
Results. Table 8 lists four benchmark contracts and shows the runtime gas overhead incurred by the application of various individual and combined invariant guards. Specifically, the OB invariant introduces the highest gas overhead among all individual invariants. This is attributable to the fact that OB utilizes a new contract state variable in storage to store its hash and, therefore, necessitates a storage load or store each time it is executed. In contrast, other invariants do not require additional storage variable access. DFU has the least impact on gas overhead, largely because it merely sets an upper bound on already accessed values, whereas other invariants typically fetch opcode results for comparison. The combined invariants do not have a gas overhead significantly higher than the individual invariants. This is because the combined invariants do not access new variables, but rather utilize the existing variables to perform additional comparisons.
Answer to RQ4: The gas overheads of these invariant guards are as low as 0% - 2.55%.
6.5. RQ5: Comparative Analysis with Other State-of-the-art (SOTA) Tools
Experiment 1: Compare with InvCon+, a SOTA invariant mining tool. InvCon+ (Liu et al., 2024), a direct follow-up work of InvCon (Liu and Li, 2022), leverages transaction pre/post conditions to generate invariants aimed at mitigating real-world smart contract vulnerabilities. InvCon only infers likely invariants and only produces raw results from Daikon (Ernst et al., 2007). In contrast, InvCon+ generates accurate invariants that are verified against the contract’s transaction history. Similar to Trace2Inv, InvCon+ takes a target contract and its transaction history as inputs and automatically generates invariants. We contacted authors of InvCon+ and obtained the InvCon+ tool. We then applied it on our benchmarks in Table 4. Following the same methodology of RQ1 and RQ3, we used $70\%$ of the transaction history for training, testing the generated invariants on the remaining $30\%$.
Experiment 2: Compare with TxSpector, a SOTA transaction attack detection tool. Though Trace2Inv is not primarily designed for transaction anomaly detection, we explored its ability to flag attack transactions. TxSpector (Zhang et al., 2020), a SOTA framework for transaction attack detection, identifies attacks using eight detectors: re-entrancy, unchecked call, failed send, timestamp dependence, unsecured balance, misuse of origin, suicidal, and gas-related re-entrancy. TxSpector takes a transaction as input and automatically flags it if it is identified as an attack. We obtained TxSpector from its public repository [^1] https://github.com/OSUSecLab/TxSpector, and applied it to the testing set of our benchmarks. Same as what is used in TxSpector paper (Zhang et al., 2020), we assigned a timeout of $60$ seconds for all benign transactions. For hack transactions which are usually more complex, we assigned a timeout of $2$ hours.
Table 9. Comparison among Trace2Inv, InvCon+, and TxSpector. (TxSpector only takes transactions as input, thus it does not have statistics about contracts.)
Trace2Inv InvCon+ TxSpector
# Contracts Applied(42) 42 27 -
# Contracts Protected(42) 31 8 -
# Hacks(27) 20 Blocked 6 Blocked 7 Detected
Average # Invariants per Contract 12 2054 -
Average FP rate(%) 0.32 73.55 15.30
Results. As shown in Table 9, InvCon+ was applied to $27$ contracts. The other $15$ contracts in our benchmarks, which utilize a proxy-implementation pattern as described in Section 3, could not be processed by InvCon+ due to its requirement for contract logic and storage to be unified. InvCon+ encountered errors for $16$ benchmarks when processing their transaction histories, and successfully generated invariants for $11$ victim contracts. These invariants secured $8$ contracts across $3$ hack incidents (Cheesebank, Punk, Warp). However, the enhanced security comes at a substantial cost: an average of $2054$ invariants per contract and a false positive rate of 73.55% if directly applying all of the invariants. This makes InvCon+ not suitable for practical application without significant human efforts to filter out unproductive invariants.
TxSpector correctly identified 7 out of 27 transactions as malicious. For the remaining 20 hack transactions, TxSpector experienced a timeout on 1 transaction and failed to flag 19 transactions. Although TxSpector reliably identifies single contract re-entrancy attacks, it struggles with detecting other attack types such as cross-contract re-entrancy and oracle manipulation. Additionally, it inaccurately marked 15.30% of benign transactions as malicious, compromising its real-world utility.
In contrast, Trace2Inv, utilizing the invariant template EOA $\land$ (OB $\lor$ DFU), effectively secured 31 out of 42 victim contracts across 20 hacks with a remarkably low false positive rate of just 0.32%. Moreover, Trace2Inv demonstrated enhanced practicality by generating an average of only $12$ invariants per contract, significantly outperforming InvCon+ in terms of real-world viability. Unlike TxSpector, which only detects hacks, Trace2Inv not only blocks a greater number of hacks but also achieves this with a significantly lower false positive rate, indicating its effectiveness in anomaly detection as well.
Identifying New Exploits: In the development of Trace2Inv, we surprisingly found two previously unreported exploit transactions, earlier than any reported exploit transactions against RariCapital1222https://etherscan.io/tx/0x4764dc6ff19a64fc1b0e57e735661f64d97bc1c44e026317be8765358d0a7392 and Yearn, [^3] https://etherscan.io/tx/0x59faab5a1911618064f1ffa1e4649d85c99cfd9f0d64dcebbc1af7d7630da98b respectively. These two transactions were initially reported by Trace2Inv as false positives, because they were not flagged as attacks when we collected benchmarks. However, after manual investigation, we found that these two transactions caused a huge financial loss and the addresses of the originators of the two transactions are flagged on EtherScan as “Rari Capital Exploiter” and “Yearn (yDai) Exploiter”, respectively. Thus, we believe these are indeed exploit transactions. This discovery further underscores Trace2Inv’s potential in unveiling new exploits.
Answer to RQ5: Trace2Inv outperforms current SOTA works on smart contract invariant mining and transaction attack detection in terms of both practicality and accuracy.
6.6. Threats to Validity
The internal threat to validity mainly lies in human mistakes in the study. Specifically, when analyzing the possibilities of bypassing invariant guards, we may miss some possible bypassing strategies. To mitigate this threat, two of the authors independently labeled the results, and whenever a conflict arises, it was resolved by the third author. All authors have more than two years’ smart contract security analysis experience.
The external threat to validity lies in the subject selection of our study. The type of hacks studied in our experiments may be limited and biased. To mitigate this issue, we systematically collected all the well-known hacks from a diverse set of sources and finally included $27$ representative hacks in our benchmark. These attacks attribute to many different root causes, including compromised keys, hash collisions, oracle manipulation, etc. Their affected contracts are from diverse application domains, e.g., bridges, lending, and yield-earning. Therefore, we believe they are representative and can be used to evaluate the effectiveness of the invariant guards.
7. Discussion
This study contributes valuable insights into the application and effectiveness of invariant-based security measures across different types of decentralized finance (DeFi) protocols. In this section, we discuss some key takeaways and their implications for future research and practical applications.
Choosing Complementary Invariants. Our findings underscore the importance of selecting a diverse set of invariants to safeguard smart contracts. Each type of invariant serves as a unique line of defense against abnormal transactions. For instance, invariants in time lock category act as temporal barriers, making it difficult for attackers to execute key functions like withdraw multiple times within a single transaction. On the other hand, invariants in data flow and money flow categories limit the token amounts that can be withdrawn in one function call. By employing a combination of these invariants, developers force attackers to only withdraw a controlled amount of tokens per transaction, which may disrupt the underlying logic that the attack relies on, thereby blocking the attack.
Dynamic Parameter Updates for Invariants. Another key insight from our research is the need for dynamic parameter updates for certain types of invariants. For invariants that are tied to variables that change cumulatively over time—like oracle prices or storage values—parameters in the invariants can quickly become obsolete. If such an invariant is violated, it could lead to a cascade of failed transactions, causing a high False Positive rate. Therefore, it is crucial for developers to continuously monitor these variables and adjust the invariant parameters. Conversely, for invariants related to independent actions like token transfers, the parameters can remain relatively stable, as user behavior in these domains tends to be stable over time.
Mitigating Flash Loan Attacks. Our benchmarks indicate a significant prevalence of flash loan-based exploits, with 17 out of 27 examined exploits leveraging flash loan. Flash loans enable users to borrow large amounts of tokens for the duration of a single transaction, providing attackers with substantial resources to execute complex exploits. Our approach of enforcing invariants can effectively mitigate the risk posed by flash loans. Many effective invariants such as EOA and OB block flash loan attacks by enforcing attackers to split their transaction into multiple ones. Without flash loan, the attacker would need to use their own assets to execute the attack with the uncertainty of other bots’ backrunning between the hackers’ transactions. This raise not only the technical but also financial barriers to successful attacks.
The Impact of Invariants on Contract Composability. Incorporating invariant guards into smart contracts might limit their adaptability and integration with DeFi protocols. However, our study in Section 4 reveals that many invariants stem from existing DeFi protocol requirements, underscoring the preference of developers for security benefits over flexibility. Moreover, our findings in RQ2, as discussed in Section 6, show that certain invariants, like OB, has almost no false positives, while others, such as TOU and DFU, can possibly block malicious transactions while allowing normal users to circumvent them. These insights imply that with careful selection and understanding of target user behaviors, developers can devise invariant guards that minimally affect contract composability.
8. Related Work
Smart Contract Invariants. There exist several works studying smart contract invariants (Liu et al., 2022a, b; Deng et al., 2024; Zhou et al., 2020). Zhou et al. (Zhou et al., 2020) introduces $6$ invariants to defend against different hacks. Cider (Liu et al., 2022a) leverages deep reinforcement learning on smart contract source code to learn invariants that prevent arithmetic overflows. SPCon (Liu et al., 2022b) utilizes function callers in past transactions of a contract to recover a likely access control model. More recently, Over (Deng et al., 2024) infers safety constraints on oracles from the contract’s source code and the history of oracle updates. Compared to the above works, Trace2Inv studies a broader range of invariants and evaluates their effectiveness and bypassability against real-world attacks.
Smart Contract Security Analysis. There are a large body of works on detection of smart contract security vulnerabilities (Feist et al., 2019; Brent et al., 2020; Tikhomirov et al., 2018; Sofware Reliability Lab, 2019; Feng et al., 2019; Luu et al., 2016; man, 2019; myt, 2019; Kalra et al., 2018; Jiang et al., 2018; Wang et al., 2020; Wüstholz and Christakis, 2020; Trail of Bits, 2019). Oyente (Luu et al., 2016) is one of the earliest symbolic execution-based security tools to detect reentrancy, mishandled exception, transaction order dependence, and timestamp dependence. It has been extended to detect greedy, prodigal and suicidal contracts (Nikolić et al., 2018). Other well-known symbolic-execution tools also include Manticore (man, 2019) and Mythril (myt, 2019) which are able to find other types of vulnerabilities, e.g., dangerous delegatecall, integer overflow, etc. Slither (sli, 2021) is another popular static security analysis tool for smart contracts. It performs data flow and control flow dependency analysis to support up to 87 bug detectors including dangerous strict equality and the dangerous usage of tx.origin. Other static analyzers include SmartCheck (Tikhomirov et al., 2018) mainly targeting bad coding practices, Securify (Sofware Reliability Lab, 2019) and Ethainter (Brent et al., 2020) for finding information-flow vulnerabilities. Moreover, many dynamic analysis tools (Jiang et al., 2018; Atzei et al., 2017; Nguyen et al., 2020; Trail of Bits, 2019; Wüstholz and Christakis, 2020; Wang et al., 2020; Liu et al., 2020) were proposed to detect smart contract vulnerabilities through fuzzing and model-based testing.
There is a large body of work on the detection and exploitation of DeFi vulnerabilities (Wu et al., 2021; Kong et al., 2023; Qin et al., 2021; Baum et al., 2022; Xue et al., 2022; Deng et al., 2023; Gudgeon et al., 2020). Qin et al. (Qin et al., 2021) and Zhou et al. (Zhou et al., 2021) formulated oracle price manipulation and arbitrage as an optimization problem to identify attack opportunities maximizing attacker’s profit. Wu et al. (Wu et al., 2021) identified several oracle price manipulation patterns from on-chain transaction data to detect real-world attacks while Kong et al. (Kong et al., 2023) detects price manipulation vulnerabilities in DeFi applications through inter-contract taint analysis. Also, Gudgeon et al. (Gudgeon et al., 2020) showcased how to explore flashloan to conduct governance attack. Baum et al. (Baum et al., 2022) surveyed the state-of-the-art mitigation techniques for front-running in DeFi, such as fair ordering, batching of blind inputs, private and secrete state. Interestingly, several works (Xue et al., 2022; Deng et al., 2023; Zhang et al., 2023) explored front-running as a defense mechanism against smart contract exploits. Different from the over-generalized security patterns used by the existing tools, our invariant guards capture the subtle semantic constraints of specific smart contracts.
Runtime Verification and Validation. Runtime verification is a light-weight verification approach that monitors the operation of a running system, detects and possibly reacts to unsatisfied observed behaviors against certain properties. Runtime verification can be used for validation purpose where the runtime checks are on properties from users’ expectation rather than from formal program specifications (Magazzeni et al., 2017). Sereum (Rodler et al., 2018) is a general runtime validation framework to protect deployed contracts against reentrancy attacks. Sereum extends Ethereum by introducing the detection module for monitoring attacks rooted in different types of reentrancy. Solythesis (Li et al., 2020) provides a source-to-source compiler that facilitates the runtime validation of smart contracts. Specifically, Solythesis takes as inputs a smart contract code and a user specified invariant. It generates an enhanced smart contracts that reject all unexpected contract transactions. Additionally, the overhead of the runtime validation approach is shown to be quite small. Trace2Inv distinguishes itself from the above-mentioned approaches in that the existing approaches only offer limited protection for several fixed types of attacks. In contrast, Trace2Inv is not bound to specific attack types and has been proven effective in mitigating many sophisticated attack vectors.
9. Conclusion
In this paper, we present the first comprehensive study of the effectiveness of practical invariant guards on preventing DeFi smart contract attacks. Our large-scale experiments on real-world DeFi hacks demonstrate that the inferred invariant guards are very effective in stopping the existing hacks, but some of the invariant guards can be bypassable by experienced attackers. We also found combining multiple invariants can be more effective than individual invariants with a lower false positive rate.
10. Data Availability
The experimental data of this paper is available on our website (authors of this paper, 2023). The artifact will be open sourced after the paper is accepted.

### v2

@@
-Demystifying Invariant Effectiveness for Securing Smart Contracts
+Demystifying Invariant Effectiveness for Securing Smart Contracts (Extended Version)
 Abstract.
@@
 Our analysis also shows that most of the invariants remain effective even when the experienced attackers attempt to bypass them.
-Additionally, we studied the possibility of combining multiple invariant guards, resulting in blocking up to $23$ of the $27$ benchmark exploits and achieving false positive rates as low as $0.32\%$.
-Trace2Inv outperforms current state-of-the-art works on smart contract invariant mining and transaction attack detection in terms of both practicality and accuracy.
-Though Trace2Inv is not primarily designed for transaction attack detection, it surprisingly found two previously unreported exploit transactions, earlier than any reported exploit transactions against the same victim contracts.
-copyright: nonecopyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXbooktitle: Companion Proceedings of the 32nd ACM Symposium on the Foundations of Software Engineering (FSE ’24), July 15–19, 2024, Porto de Galinhas, Brazilprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
+Additionally, we studied the possibility of combining multiple invariant guards, resulting in blocking up to $23$ of the $27$ benchmark exploits and achieving false positive rates as low as $0.28\%$.
+Trace2Inv significantly outperforms state-of-the-art works on smart contract invariant mining and transaction attack detection in accuracy.
+Trace2Inv also surprisingly found two previously unreported exploit transactions.
+runtime validation, invariant generation, dynamic analysis, flash loan
+copyright: rightsretaineddoi: 10.1145/3660786journalyear: 2024submissionid: fse24main-p241-pjournal: PACMSEjournalvolume: 1journalnumber: FSEarticle: 79publicationmonth: 7ccs: Security and privacy Software security engineeringccs: Software and its engineering Software testing and debugging
 1. Introduction
 Blockchain technology has paved the way for decentralized, resilient, and programmable ledgers on a global scale.
-One of its most impactful applications is smart contracts, which can be deployed onto a blockchain.
+One of its most impactful applications is smart contracts.
 These smart contracts allow developers to encode intricate transaction rules that govern the ledger.
 This innovation has made both blockchains and smart contracts essential infrastructure for decentralized financial services, commonly known as DeFi.
-As of Sept 25, 2023, the Total Value Locked (TVL) in $2,933$ DeFi protocols has reached an impressive $48.58$ billion (def, 2023a).
+As of Sept 25th, 2023, the Total digital asset Value Locked (TVL) in $2,933$ DeFi protocols has reached an impressive $48.58$ billion (def, 2023a).
 However, the landscape is not without its challenges.
 Security attacks pose a significant threat to the security of smart contracts.
-Attackers can exploit various vulnerabilities by sending malicious transactions, potentially leading to the theft of millions of dollars from users.
-As of Sept 25, 2023, the financial losses attributed to security attacks on DeFi protocols exceeded $5.53$ billion USD (def, 2023b).
+Attackers can exploit various vulnerabilities by sending malicious transactions, potentially leading to the theft of millions of dollars.
+As of Sept 25th, 2023, the financial losses attributed to security attacks on DeFi protocols exceeded $5.53$ billion USD (def, 2023b).
 One key observation is that transactions initiated by attackers often display abnormal behaviors when compared to standard transactions from regular DeFi contract users.
@@
 These checks involve enforcing various runtime invariants, such as restricting the maximum number of digital asset deposits or withdrawals in a contract to prevent market manipulation.
-Another example is to limit the interaction of other contracts to prevent attackers from crafting sophisticated attack strategies.
-However, these mechanisms are often manually designed and tailored for specific contracts.
+Another example is to limit the interaction between contracts to prevent attackers from crafting sophisticated attack strategies.
+However, these mechanisms are often manually designed and tailored for specific DeFi protocols.
 This raises questions about their effectiveness across different types of contracts and whether they maintain an acceptable false positive rate without hindering normal user activities.
@@
 Our findings indicate that dynamically inferred dynamic invariants serve as effective mechanisms for thwarting security breaches.
-When appropriately configured, these invariants can neutralize over 74.1% of malicious attacks while maintaining a false positive rate of less than 0.32%.
+This paper then proposes new strategies to combine multiple invariants effectively.
+Our combined invariants can neutralize over 74.1% of malicious attacks while maintaining a false positive rate of less than 0.28%.
 Trace2Inv: To facilitate this study, we have developed Trace2Inv, a scalable and extensible invariant synthesis framework.
@@
 2. Background
-Blockchain is a distributed, immutable ledger technology that records transactions across multiple nodes in a network.
+Blockchain is a distributed and immutable ledger technology that records transactions across multiple nodes in a network.
 It employs cryptographic techniques to ensure data integrity and consensus algorithms to maintain final consistency across all participating nodes.
-Smart contracts are self-executing contracts with the terms of the agreement directly written into code.
+Smart contracts are self-executing programs with the terms of the agreement directly written into code.
 Deployed on blockchains, they are immutable and transparent, enabling trustless transactions without the need for intermediaries.
-Invariant guards (also called circuit breakers) are runtime checks around contract invariant conditions that shall always hold during contract exeuction, aiming to secure smart contracts on the fly.
+Invariant guards (also called circuit breakers) are runtime checks around contract invariant conditions that shall always hold during contract execution, aiming to secure smart contracts on the fly.
 Ethereum Virtual Machine (EVM) is the runtime environment for smart contract execution on Ethereum.
-It is a Turing-complete virtual machine that interprets and executes the bytecode compiled from contracts programmed in high-level language like Solidity.
+It is a Turing-complete virtual machine that interprets and executes the bytecode compiled from contracts programmed in a high-level language like Solidity.
 Gas is a unit of transaction fee on Ethereum, used to quantify the computational efforts for the execution of EVM operations.
@@
 They are more subtle to detect and attacks usually involve more sophisticated steps (Zhou et al., 2023).
-DeFi protocols are major targets for smart contract attacks, which experienced \$5.53 billion loss out of the overall \$6.94 billion in recent blockchain incidents (def, 2023b).
+DeFi protocols are major targets for smart contract attacks, which have experienced \$5.53 billion loss out of the overall \$6.94 billion caused by recent blockchain incidents (def, 2023b).
 3. Motivating Example
@@
 In the exploit transaction, the attacker distorts the market of Curve to cause the Harvest Finance to make sub-optimal investment decisions.
-Specifically, the attacker in the exploit transaction first borrows a large amount of digital assets and uses the borrowed asset to buy USDC in Curve to inflate the USDC price in Curve.
+Specifically, the attack transaction first borrows a large amount of digital assets and uses the borrowed asset to buy USDC in Curve to inflate its price.
 Then it deposits $49.98$M USDC into Harvest vault contract, which increases its USDC balance from $72.83$M to $122.51$M.
@@
 Remarkably, this identical attack vector is executed three times within one exploit transaction, consuming an unusually high gas count of $9,895,111$, narrowly within the gas limit of $12,065,986$ at the time.
+More details of this example can be found at the website (Chen et al., 2024e).
 Abnormal Behaviors: We identify four distinct dimensions of abnormal behavior: a high frequency of user interactions with the Harvest vault contract, an exceptionally large volume of token flow, abrupt fluctuations in the total supply of fUSDC tokens, and remarkably high gas consumption.
@@
 Specifically, we have observed that the exploit transaction is the first in the contract’s history to: (1) invoke the withdraw function from a contract rather than from a user address, (2) call both deposit and withdraw functions $3$ times within one transaction, (3) consume more gas than any previous transaction, (4) withdraw more USDC from the protocol than any other transaction, (5) elevate the total supply of fUSDC tokens to an all-time high.
+We observed several other outlier data points besides the last data point.
+For access frequency, two blocks contained two withdraw invocations within the same block, which, upon investigation, were found to originate from separate addresses.
+We believe it was a coincidence that two users withdrew their funds in the same block.
+For token flow, a very large USDC deposit occurred immediately after the contract’s deployment, which, upon manual analysis, was identified as an initial liquidity provided by the Harvest Finance team.
+Similarly, the abrupt jump in the total supply of fUSDC tokens in the early stage was also due to this initial liquidity provision.
+Since users receive fUSDC tokens when they deposit USDC into the vault and these tokens are burned upon withdrawal of USDC, a large initial deposit would lead to a corresponding increase in the total supply of fUSDC tokens.
+For gas consumption, a transaction at block 11094671 consumed 5,062,934 gas, which, after manual analysis, was found to be sent by the Harvest Finance team for migrating old vaults to new ones.
+This migration process involved numerous operations in a single transaction to ensure atomicity, leading to high gas consumption.
+Overall, while there were anomalies during the early stage of the contract’s deployment, the different transaction metrics stabilized after some blocks until significantly changed by the exploit transaction.
+If contract maintainers can identify the ”stable” stage and use these transactions to infer invariants, the inferred invariants can be much more accurate.
+However, in this paper, we assume such information is not available.
 Apply Inferred Invariants: The multi-dimensional abnormalities observed in the exploit transaction highlight a stark departure from typical transactional behaviors.
@@
 RQ4: Invariant guards require additional gas at runtime. What are the gas overheads of different types of invariant guards?
-RQ5: In terms of enhancing smart contract security, how does this work compare to other state-of-the-art works in smart contract invariant generation or transaction anomaly detection?
+RQ5: In terms of enhancing smart contract security, how does this work compare to other state-of-the-art works in contract invariant generation or transaction anomaly detection?
 4. Invariants
@@
 Additionally, we conducted a qualitative study on both audit reports and source code of the $63$ audited projects from ConsenSys, a leading smart contract auditing firm, from May 2020 to March 2023 (con, 2023).
-One author extracted $2181$ enforced invariants from the audit reports and smart contracts under audit by searching for keywords such as “require” and “assert”, after eliminating any duplicates.
+One author extracted $2,181$ enforced invariants from the audit reports and smart contracts under audit by searching for keywords such as “require” and “assert”, after eliminating any duplicates.
 The author then manually reviewed these invariants to extract templates for pattern matching against remaining uncategorized invariants.
@@
 Table 1.
-Qualitative Study Statistics Overview (The table’s left section presents key statistics from the qualitative study.
+Qualitative Study Statistics Overview+ (The table’s left section presents key statistics from the qualitative study.
 The middle section presents categorized instances across invariant categories.
@@
 Data Flow 181
-The above process resulted in $826$ invariants under $8$ categories, which comprise $37.87\%$ of the overall $2181$ invariants, as shown in Table 1.
-Access Control and DataFlow are the top and second common categories, respectively.
-The remaining invariants ($62.13\%$) can not be categorized for various reasons.
+The above process resulted in $826$ invariants under $8$ categories, which represent $37.87\%$ of the $2181$ invariants, as shown in Table 1.
+Access Control and Data Flow are the two common categories.
+The remaining invariants ($62.13\%$) were not categorized for various reasons.
 The most common reason is that invariants are specific to a particular protocol.
@@
 Category ID Name Template Parameter References
-Access Control EOA onlyEOA $\mathsf{msg.sender}=\mathsf{tx.origin}$ (Liu et al., 2022b; Ghaleb et al., 2023)
+Access Control EOA onlyEOA $\mathsf{msg.sender}=\mathsf{tx.origin}$ (Liu et al., 2022b) (Ghaleb et al., 2023)
 SO isSenderOwner $\mathsf{msg.sender}=\mathsf{owner?}$ address
@@
 Oracle Slippage OR OracleRange $\mathsf{pr_{LB}?}\leq\mathsf{r_{newPr}}\leq\mathsf{pr_{UB}?}$ Integer (dfo, 2023b)
-OD OracleDeviation $|(\mathsf{r_{newPr}}-\mathsf{x_{oldPr}})/\mathsf{x_{oldPr}}|\leq\mathsf{pr_{Dev}?}$ Integer (dfo, 2023b; Breidenbach et al., 2021)
-Special Storage TSU TotalSupplyUpperBound $\mathsf{x_{totSup}}\leq\mathsf{totSup?}$ Integer (Aav, 2023; dfo, 2023a)
-TBU TotalBorrowUpperBound $\mathsf{x_{totBor}}\leq\mathsf{totBor?}$ Integer (Aav, 2023; dfo, 2023a)
-MoneyFlow TIU TokenInUpperBound $\mathsf{r_{tokenIn}}\leq\mathsf{v?}$ Integer (bal, 2023; dfo, 2023a)
+OD OracleDeviation $|(\mathsf{r_{newPr}}-\mathsf{x_{oldPr}})/\mathsf{x_{oldPr}}|\leq\mathsf{pr_{Dev}?}$ Integer
+(dfo, 2023b)
+(Breidenbach et al., 2021)
+Special Storage TSU TotalSupplyUpperBound $\mathsf{x_{totSup}}\leq\mathsf{totSup?}$ Integer (Aav, 2023) (dfo, 2023a)
+TBU TotalBorrowUpperBound $\mathsf{x_{totBor}}\leq\mathsf{totBor?}$ Integer (Aav, 2023) (dfo, 2023a)
+Money Flow TIU TokenInUpperBound $\mathsf{r_{tokenIn}}\leq\mathsf{v?}$ Integer (bal, 2023) (dfo, 2023a)
 TIRU TokenInRatioUpperBound $\mathsf{r_{tokenIn}}\leq\mathsf{v?}$ Integer (bal, 2023)
-TOU TokenOutUpperBound $\mathsf{r_{tokenOut}}/\mathsf{b_{token,adr}}\leq\mathsf{v?}$ Integer (bal, 2023; dfo, 2023a)
+TOU TokenOutUpperBound $\mathsf{r_{tokenOut}}/\mathsf{b_{token,adr}}\leq\mathsf{v?}$ Integer (bal, 2023) (dfo, 2023a)
 TORU TokenOutRatioUpperBound $\mathsf{r_{tokenOut}}/\mathsf{b_{token,adr}}\leq\mathsf{v?}$ Integer (bal, 2023)
@@
 DFL DataFlowLowerBound $\mathsf{var?}\geq\mathsf{v?}$ Integer
-4.1. Access Control
+4.1. Access Control (also called Permission Control)
 Many research papers have conducted extensive studies on the access control (Liu et al., 2022b; Ghaleb et al., 2023).
@@
 For example, if user address $a$ calls contract $b$ which in turn calls contract $c$, during the execution of $c$, the sender address is $b$ while the origin address is $a$.
-onlyEOA (EOA): This template verifies that the transaction’s origin matches the sender’s address, thereby confirming it was initiated from an externaly owned user address (i.e., EOA address) rather than a contract address.
+onlyEOA (EOA) [^1] It is important to note that the forthcoming Spectra upgrade of Ethereum, anticipated for late 2024 or early 2025, will implement EIP-3074 (EIP, 2024b).
+This implementation is expected to make EOA completely bypassable.
+This template verifies that the transaction’s origin matches the sender’s address, thereby confirming it was initiated from an externaly owned user address (i.e., EOA address) rather than a contract address.
 The intuition of this invariant template is that many attack strategies involves multiple sophisticated interactions and therefore attackers often have to write their own contracts.
@@
 4.5. Oracle Slippage
-The Oracle Slippage category mitigates risks tied to price oracles in decentralized finance (DeFi) applications.
+The Oracle Slippage category mitigates risks tied to price oracles in DeFi applications.
 The intuition of templates in this category is to detect potential price manipulation by malicious attacks.
@@
 The Data Flow category is subdivided into four specific invariants, each designed to address a particular type of variables in the data flow.
-MappingUpperBound (MU): This invariant focuses on values stored in the contract’s mapping data structure, often representing a user’s property(e.g., a user’s share in a liquidity pool).
+MappingUpperBound (MU): This invariant focuses on values stored in the contract’s mapping data structure, often representing a user’s property(e.g., a user’s shares/balances).
 To constrain such user-specific values, we introduce a parameter $\mathsf{v?}$ to set an upper limit on these mappings.
@@
 This is often in functions that initiate token transfers.
-These invariants helps the contract ensure that every value used for the calculation of token transfers is within the normal range.
+These invariants help the contract ensure that every value used for the calculation of token transfers is within the normal range.
 5. Trace2Inv
-We present Trace2Inv, our proposed framework to infer concrete invariants as introduced in Section 4 for a given contract by analyzing its historical transaction traces.
-As shown in Figure 2, Trace2Inv is structured into three modules: trace parser, invariant-related data extraction, and invariant generation.
-The second module consists of three sub-modules: invocation tree analysis, type inference, and dynamic taint analysis.
+We present Trace2Inv, shown in Figure 2, a framework to infer concrete invariants as introduced in Section 4 for a contract by analyzing its transaction traces.
+Trace2Inv consists of three modules: trace parser, invariant-related data extraction, and invariant generation.
+The second module consists of three submodules: invocation tree analysis, type inference, and dynamic taint analysis.
 Figure 2. An Overview of Trace2Inv
@@
 5.2. Invocation Tree Analysis
-Invocation tree analysis extracts invariant-related data exclusively from the invocation tree, sufficient to collect data for access control, time lock, gas control, re-entrancy, oracle, and money flow invariants.
-For instance, for the sender opcode, its results can be extracted from the invocation tree by searching for parent node of the node of target contract.
-Moreover, the invocation tree can also be used to identify locations of re-entrancy by capturing the nested and recursive calls to the target contract.
+Invocation tree analysis extracts invariant-related data from the invocation tree, that is sufficient to collect data for access control, time lock, gas control, re-entrancy, oracle, and money flow invariants.
+For instance, for the sender opcode, its results is extracted from the invocation tree by searching for parent node of the target contract node.
+Moreover, the invocation tree is also used to identify locations of re-entrancy by capturing nested and recursive calls to the target contract.
 Oracle values are also obtained by traversing the invocation tree to locate calls to the oracle.
-For money flow invariants, the required values can also be directly read from function calls of transfering Ether/ERC20.
+For money flow invariants, the required values are also read from function calls of transferring Ether/ERC20.
 5.3. Type Inference
 Type inference in Trace2Inv infers the type of storage slot when it is accessed.
-It is essential for extracting data relevant to both special storage and data flow invariants, as accurately decoding storage accesses with the correct type is necessary for these invariant categories.
+It is essential for extracting data relevant to both special storage and data flow invariants, as accurately decoding storage accesses with the correct type is necessary for those invariants.
 Decoding storage slots is straightforward when they are listed in the contract’s storage layout, which is typically the case for special storage invariants.
 However, the challenge arises with complex data structures like mapping, often used in data flow invariants.
-These structures may reference storage slots not explicitly present in the contract’s storage layout, which instead are computed through operations such as sha3 and arithmetic functions.
-To solve this issue, Trace2Inv maintain a preimage dictionary to track the key’s computation.
-When Trace2Inv encounters a sha3 opcode, the mapping positions and slots are recorded.
+These structures may reference storage slots not explicitly present in the contract’s storage layout, which are instead computed through operations such as sha3 and arithmetic functions.
+To solve this issue, Trace2Inv maintains a preimage dictionary to track the key’s computation.
+When we encounter a sha3 opcode, the mapping positions and slots are recorded.
 In Solidity, the first $32$-bytes represent the mapping slot, and the next $32$-bytes serve as the mapping position.
 In Vyper, these roles are reversed.
-Anytime a 64-byte sha3 hash is encountered, both its hash value and origin are recorded for future reference.
-When an sload operation has a key that is not present in the storage layout, the preimage dictionary is consulted.
+Anytime a 64-byte sha3 hash is encountered, both its hash value and origin are recorded.
+When a sload operation has a key that is not present in the storage layout, the preimage dictionary is consulted.
 We recursively trace back the computation steps until we identify a mapping data structure that is in the storage layout.
-Utilizing the types of the mapping data structure, we can infer the type of storage slot accessed.
+Using the types of the mapping data structure, we infer the type of storage slot accessed.
 5.4. EVM-level Dynamic Taint Analysis
@@
 Category Opcodes and Locations
-Sources External Address Variables balance, extcodesize, extcodecopy, extcodehash,
-Execution Context Variables origin, caller, address, codesize, selfbalance, pc, msize, gas
-Call Data Variables callvalue, calldataload, calldatasize, calldatacopy
-Return Data Variables returndatasize, returndatacopy
-Block Variables blockhash, coinbase, timestamp, number, prevrandao, gasprice, gaslimit, chainid,
+Sources External Address balance, extcodesize, extcodecopy, extcodehash
+Execution Context origin, caller, address, codesize, selfbalance, pc, msize, gas
+Call Data callvalue, calldataload, calldatasize, calldatacopy
+Return Data returndatasize, returndatacopy
+Block blockhash, coinbase, timestamp, number, prevrandao, gasprice, gaslimit, chainid
 Storage sload(untainted)
@@
 Bit Level Taint Propagation.
-Our taint analyzer maintains three distinct taint trackers: the stack tracker, the memory tracker, and the storage tracker.
+Our taint analyzer maintains three distinct taint trackers: the stack, the memory, and the storage trackers.
 Initially, all these trackers are empty.
@@
 5.6. Trace2Inv Implementation
-Trace2Inv’s input, transaction trace data, can be obtained from Ethereum via its API debug_traceTransaction.
+Trace2Inv’s input, transaction trace data, can be obtained from Ethereum via its API debug_trace-Transaction.
 Within the Trace Parser module, Trace2Inv queries EtherScan (Eth, 2020) to gather additional transaction metadata not directly available in the trace data, such as block number and transaction origin.
@@
 In such cases, Trace2Inv fetches the storage layout from EVM Storage (EVM, 2024).
-Several optimizations are incorporated into Trace2Inv implementation to enhance its performance.
-Firstly, when fetching trace data from an Ethereum archive node, we optimize the process by using batching RPC requests.
+Several optimizations are incorporated into Trace2Inv to enhance its performance.
+First, when fetching trace data from an Ethereum archive node, we optimize the process by using batching RPC requests.
 This significantly reduces the overhead associated with individual API calls.
-Secondly, we employ parallelization techniques to speed up the process of fetching and parsing trace data.
-Specifically, multi-processing is used to concurrently handle different segments of trace data, converting them into summaries in a more time-efficient manner.
-Thirdly, caching is utilized to further optimize performance; the system caches query results from EtherScan, archive nodes, and compilers.
+Second, we use parallelization techniques to speed up the fetching and parsing trace data.
+Specifically, multiprocessing is used to concurrently handle different segments of trace data, converting them into summaries in a time-efficient manner.
+Third, caching is utilized to further optimize performance; the system caches query results from EtherScan, archive nodes, and compilers.
 This minimizes redundant executions and API calls, thereby accelerating the overall analysis process.
 6. Evaluation
+In this section, we aim to empirically answer the research questions raised in Section 3 by applying invariant guards to real-world exploits on Ethereum Blockchain.
+We systematically collected economic exploit incidents that cost greater than 300K USD financial loss from February 14, 2020, to August 1, 2022 on Ethereum Blockchain.
+Our benchmark is compiled from a diverse set of sources including academic publications (Chen et al., 2022; Qin et al., 2021), industry databases (SlowMist, 2023; BlockSec, 2023), and open-source GitHub repositories (Contributors, 2023a, b).
+It is worth noting that we exclude from our benchmarks any hacks targeting individual user wallets, as these are primarily the result of private key leakage, rather than protocol vulnerabilities.
+We also exclude hacks where the victim contracts are close-source, as our manual analysis requires the source code of the victim contracts.
+Note that Trace2Inv can also be applied to close-source contracts, as long as their function ABIs and storage layout are available.
 Table 4.
-Benchmarks.
+Benchmarks+.
 (Exploit: the first exploit transaction during the incident.
@@
     Kind Victim Protocol Root Cause Date Loss Exploit FL Contracts Type PL History
-    Bridges RoninNetwork keys compromised 22/03/29 624M (Etherscan, 2023u) RoninNetwork I+P S 95345
-HarmonyBridge keys compromised 22/06/24 100M (Etherscan, 2023h) HarmonyBridge P S 32149
+    Bridges RoninNetwork keys compromised 22/03/29 624M (Ron, 2023) RoninNetwork I+P S 95345
+HarmonyBridge keys compromised 22/06/24 100M (Har, 2023a) HarmonyBridge P S 32149
 Nomad
@@
 a valid root
-22/08/01 152M (Etherscan, 2023m) Nomad I+P S 15630
-PolyNetwork hash collision 21/08/10 611M (Etherscan, 2023p) PolyNetwork I+P S 44509
-    Lending bZx2 oracle manipulation 20/02/18 630K (Etherscan, 2023b) bZx2 I+P S 711
+22/08/01 152M (Nom, 2023) Nomad I+P S 15630
+PolyNetwork hash collision 21/08/10 611M (Pol, 2023) PolyNetwork I+P S 44509
+    Lending bZx2 oracle manipulation 20/02/18 630K (bZx, 2023) bZx2 I+P S 711
 Warp oracle manipulation 20/12/17 8M Warp P S 148
-(Etherscan, 2023y) Warp_I I S 31
+(War, 2023) Warp_I I S 31
 CheeseBank oracle manipulation 20/11/06 3.3M CheeseBank_1 I+P S 615
 CheeseBank_2 I+P S 593
-(Etherscan, 2023c) CheeseBank_3 I+P S 557
-InverseFi oracle manipulation 22/06/16 1.26M (Etherscan, 2023l) InverseFi I+P S 7590
+(Che, 2023) CheeseBank_3 I+P S 557
+InverseFi oracle manipulation 22/06/16 1.26M (Inv, 2023) InverseFi I+P S 7590
 CreamFi1 cross contract re-entrancy 21/08/30 18M CreamFi1_1 I+P S 5
-(Etherscan, 2023d) CreamFi1_2 I+P S 58184
+(Cre, 2023a) CreamFi1_2 I+P S 58184
 CreamFi2 oracle manipulation 21/10/27 130M CreamFi2_1 I+P S 1270
@@
 CreamFi2_3 I+P S 261
-(Etherscan, 2023e) CreamFi2_4 I+P S 98
-RariCapital1 read-only re-entrancy 21/05/09 10M (Etherscan, 2023r) RariCapital1 I+P S 667
+(Cre, 2023b) CreamFi2_4 I+P S 98
+RariCapital1 read-only re-entrancy 21/05/09 10M (Rar, 2023a) RariCapital1 I+P S 667
 RariCapital2 cross contract re-entrancy 22/04/30 80M RariCapital2_1 I+P S 614
@@
 RariCapital2_3 I+P S 404
-(Etherscan, 2023s) RariCapital2_4 I+P S 776
-XCarnival logic error 22/06/26 3.87M (Etherscan, 2023z) XCarnival I+P S 342
-    Yield-Earning Harvest1 oracle manipulation 20/10/26 33.8M (Etherscan, 2023i) Harvest1 I+P S 2050
-Harvest2 oracle manipulation 20/10/26 (Etherscan, 2023j) Harvest2 I+P S 2161
-ValueDeFi oracle manipulation 20/11/14 7.4M (Etherscan, 2023w) ValueDeFi I+P S 295
+(Rar, 2023b) RariCapital2_4 I+P S 776
+XCarnival logic error 22/06/26 3.87M (XCa, 2023) XCarnival I+P S 342
+    Yield-Earning Harvest1 oracle manipulation 20/10/26 33.8M (Har, 2023b) Harvest1 I+P S 2050
+Harvest2 oracle manipulation 20/10/26 (Har, 2023c) Harvest2 I+P S 2161
+ValueDeFi oracle manipulation 20/11/14 7.4M (Val, 2023) ValueDeFi I+P S 295
 Yearn forced investment 21/02/04 11M Yearn P S 672
-(Etherscan, 2024) Yearn_I I S 26688
-VisorFi re-entrancy 21/12/21 8.2M (Etherscan, 2023x) VisorFi I+P S 1693
-UmbrellaNetwork underflow bug 22/03/20 700K (Etherscan, 2023v) UmbrellaNetwork I+P S 59
-PickleFi access control 20/11/21 20M (Etherscan, 2023o) PickleFi I+P S 5439
-    Others Eminence logic error 20/09/29 7M (Etherscan, 2023g) Eminence I+P S 20589
-Opyn logic error 20/08/04 371k (Etherscan, 2023n) Opyn I+P S 67
-IndexFi logic error 21/10/15 16M (Etherscan, 2023k) IndexFi I+P S 20641
+(Yea, 2024) Yearn_I I S 26688
+VisorFi re-entrancy 21/12/21 8.2M (Vis, 2023) VisorFi I+P S 1693
+UmbrellaNetwork underflow bug 22/03/20 700K (Umb, 2023) UmbrellaNetwork I+P S 59
+PickleFi access control 20/11/21 20M (Pic, 2023) PickleFi I+P S 5439
+    Others Eminence logic error 20/09/29 7M (Emi, 2023) Eminence I+P S 20589
+Opyn logic error 20/08/04 371k (Opy, 2023) Opyn I+P S 67
+IndexFi logic error 21/10/15 16M (Ind, 2023) IndexFi I+P S 20641
 RevestFi re-entrancy 22/03/27 11.2M RevestFi P S 1635
-(Etherscan, 2023t) RevestFi_I I S 1463
-DODO access control 21/03/08 700K (Etherscan, 2023f) DODO I+P S 42
+(Rev, 2023) RevestFi_I I S 1463
+DODO access control 21/03/08 700K (DOD, 2023) DODO I+P S 42
 Punk access control 21/08/10 8.9M Punk_1 I+P S 28
 Punk_2 I+P S 42
-(Etherscan, 2023q) Punk_3 I+P S 37
+(Pun, 2023) Punk_3 I+P S 37
 BeanstalkFarms flashloan assisted commit 22/04/16 182M BeanstalkFarms P V 5785
-(Etherscan, 2023a) BeanstalkFarms_I I S 306
-In this section, we aim to empirically answer the research questions raised in Section 3 by applying invariant guards to real-world exploits on Ethereum Blockchain.
-We systematically collected economic exploit incidents that cost greater than 300K USD financial loss from February 14, 2020, to August 1, 2022 on Ethereum Blockchain.
-Our benchmark is compiled from a diverse set of sources including academic publications (Chen et al., 2022; Qin et al., 2021), industry databases (SlowMist, 2023; BlockSec, 2023), and open-source GitHub repositories (Contributors, 2023a, b).
-It is worth noting that we exclude from our benchmarks any hacks targeting individual user wallets, as these are primarily the result of private key leakage, rather than protocol vulnerabilities.
-We also exclude hacks where the victim contracts are close-source, as our manual analysis requires the source code of the victim contracts.
-Note that Trace2Inv can also be applied to close-source contracts, as long as their function ABIs and storage layout are available.
+(Bea, 2023) BeanstalkFarms_I I S 306
 Table 4 presents the benchmark dataset in our study.
-It comprises 27 hacks which cumulatively resulted in financial losses exceeding 2 billion USD.
+It comprises 27 hacks which resulted in financial losses over $2$ billion USD.
 The column FL denotes whether the exploit involves flash loans, which require atomicity for the hack transactions.
 The column Type denotes the type of victim contracts.
-For each exploit, two types of victim contracts are manually identified: payload (P) refers to the protocol contract that eventually transfers abnormal amounts of tokens out of the protocol, and interface (I) refers to the contract that is directly invoked by users to initiate this abnormal transfer.
-The History column signifies the length of the transaction history up to the exploit transaction for each victim contract.
+For each hack, two types of victim contracts are manually identified: payload (P) refers to the contract that eventually transfers abnormal amounts of tokens out of the protocol, and interface (I) refers to the contract that is directly invoked by users to initiate this transfer.
+The History column gives the length of the transaction history up to the hack transaction for each victim contract.
 6.1. RQ1: Effectiveness of Smart Contract Invariants
@@
 On the other hand, any non-exploit transactions blocked are counted as False Positives.
+Table 5.
+Complete Results of Invariant Effectiveness Evaluation.
+(Gray cells indicate the exploit transaction is blocked by corresponding invariant guard(True Positives).
+The number in the cell represents the number of non-exploit transactions blocked by corresponding invariant guard(False Positives).
+Symbol - represents that the invariant is by nature not applicable to the benchmark contract.
+Symbol represents that the transactions in the training set violate the invariant.
+Symbol $\emptyset$ represents the transactions in the training set do not have enough data points to learn the invariant.)
+Access Control Time Lock Gas Control Re Oracle Storage Money Flow Data Flow     Kind Benchmarks     EOA     SO     SM     OO     OM     SB     OB     LU     GS     GC     RE     OR     OD     TSU     TBU     TIU     TOU     TIRU     TORU     MU     CVU     DFU     DFL RoninNetwork 0 0 0 0 - - - - 0 0 0.4 0.1 - 0 0 0 HarmonyBridge 0 0 0 0 0 2.8 2.8 0 - - - - 0 0 0 0 - 0 0 0 Nomad 0 0 0 0 0 0 0 - - - - 0 0 0 0 - - 0 0     Bridges PolyNetwork 0 0 0 0 0 0 0 0 - - - - 0 0 0 0 - 0 - - bZx2 0 0 2.3 0 2.3 0 0 3.3 1.4 1.9 0 $\emptyset$ $\emptyset$ 0.5 0 0 0 - - 0 - 0 0.5 Warp 0 0 0 2.3 0 0 0 6.8 6.8 0 - - - 93.2 0 4.5 0 0 - - 4.5 0 Warp_I 0 0 11.1 0 33.3 - - 0 22.2 33.3 - 88.9 44.4 - - - - - - - - - - CheeseBank_1 0 0 0 0 0 0.5 0.5 4.3 0 7.1 1.6 0 33.2 2.2 1.1 2.2 1.6 - - 3.3 0.5 CheeseBank_2 0 0 0 0 0 0.6 1.1 1.1 0 10.2 2.8 0 7.9 0 0 0 0 - - 0 0 CheeseBank_3 0 0 0 0 0 0 1.8 1.8 0 10.8 1.8 0 7.8 3 0 3 0 - - 3 3.6 InverseFi 0 0 0 0.1 0.1 0 0 0.3 0.1 0 0 0 0 1.9 16.3 1.5 0 1.3 0.1 - - 2 0.2 CreamFi1_1 0 0 $\emptyset$ 0 $\emptyset$ 0 0 $\emptyset$ $\emptyset$ $\emptyset$ 0 - - - - - - - - - - $\emptyset$ $\emptyset$ CreamFi1_2 0.5 0.5 0.1 0.5 0 0.5 0.1 0.3 0 - - 0 0 0 0 0.1 0.2 - 0 0 0 CreamFi2_1 0 0.3 0 0.3 0.3 0 0 0.8 1.1 1.1 0 0 0 2.9 9.5 1.8 0 0 0.8 2.6 - 0.8 0.5 CreamFi2_2 0 0 1.5 0 1.9 0 0 3.7 3.3 2.6 0 0 0 0 0 0 0 0 0.4 2.6 - 0 1.1 CreamFi2_3 3.8 0 9 0 12.8 0 0 7.7 3.8 2.6 0 - - 10.3 0 3.8 0 3.8 0 0 - 3.8 1.3 CreamFi2_4 0 0 0 0 0 0 0 0 0 0 0 - 20.7 0 3.4 0 3.4 3.4 - 3.4 0 RariCapital1 0 0 6 0 4.5 0 0 6.5 1.5 0 0 - - - - 0.5 0 - - - 0.5 0.5 0 RariCapital2_1 0 0 0 0 0 0 0 1.1 1.1 0.5 0 - - 10.9 6 0.5 0.5 0 0 - 0.5 0.5 0 RariCapital2_2 1.8 0 0.4 0 0.4 0 0 1.8 0.4 0.4 0 - - 6.7 2.2 0.9 0 0.9 0.4 1.3 - 0.9 2.7 RariCapital2_3 0 0 0 0 0 0 0 0 0.8 1.7 0 - - 0 0 0 0 0.8 0 1.7 - 0 0 RariCapital2_4 0 0 0 0 0 0 0 0.4 0.4 2.6 0 - - 0 0 0 0 0 0 0.9 - 0 0     Lending XCarnival 0 6.9 5.9 0 0 2 11.8 8.8 0 - - 0 4.9 3.9 1 4.9 28.4 3.9 3.9 1 1 Harvest1 0 0.5 0 0 0 0 0 0.2 0 0 0 29.3 0 1.1 - 0 0.2 0 0.2 - - 2.6 0.2 Harvest2 0 0 0 0 0 0 0 0 0 0 0 34.7 0 0 - 0 0 0.2 0.2 - - 1.2 0.2 ValueDeFi 0 0 0 0 0 3.4 1.1 2.3 0 64.8 0 73.9 - 0 0 0 0 - - 8 1.1 Yearn1 0 0 0 0 0 0 0 0 0 0 - - - - - - - - - - $\emptyset$ $\emptyset$ Yearn1_I 0 0 0 0 0.1 0.1 0 - - - - 0 0.1 0.1 1.8 0 - 4.4 0 VisorFi 0 0 0 0 0 0 0 0 0 - - - - 0 1.2 0 0 - - 4.2 0 UmbrellaNetwork 0 11.8 11.8 0 0 5.9 0 0 0 - - - - 5.9 0 5.9 0 - - 5.9 17.6     Yield-Earning PickleFi 0 0 0 0 0 0 0 0 0.1 0.1 0 - - - - 0 0.1 - - - - 0.1 0.1 Eminence 0 0 0 0 0 0.1 0 0 - - 20.5 - 0 0.5 0 0.1 - - 0.4 0 Opyn 0 0 0 0 0 0 0 0 0 0 0 - - - - - - - - - - $\emptyset$ $\emptyset$ IndexFi 0 0 0 0 0 0 0 0 0.7 0.1 0 - - 6.6 - 0 0 0.1 0 0.4 - 0 0 RevestFi 0 0 0 3.7 0 0 4.1 0.4 1.8 0 - - - - - - - - - - $\emptyset$ $\emptyset$ RevestFi_I 1.4 0 0 0 4.1 0 0 4.6 0.5 2.1 0 - - - - - - - - - - - - DODO 0 0 8.3 0 0 0 0 33.3 25 0 - - 0 - - - - - - - $\emptyset$ $\emptyset$ Punk_1 0 0 $\emptyset$ 0 $\emptyset$ - - 25 0 0 0 - - - - 0 0 - - - - 0 0 Punk_2 0 0 $\emptyset$ 0 0 - - 0 0 0 - - - - 0 0 - - - - 0 0 Punk_3 0 0 $\emptyset$ 0 $\emptyset$ - - 18.2 0 0 0 - - - - 0 0 - - - - 0 0 BeanstalkFarms 0 0 0.2 0 0.4 0 0.2 0.7 0.6 0 - - 31.8 - 0 0.1 0 0 - - 0.8 0.3     Others BeanstalkFarms_I 0 2.2 0 2.2 0 - - 0 0 0 - - - - - - - - - - - - - # Contracts Applied 42 39 22 39 25 33 33 37 41 41 40 11 11 21 16 34 34 28 28 11 7 33 33 # Contracts Protected 23 6 7 5 9 11 13 12 30 23 2 7 6 8 12 9 22 5 22 1 2 23 1 # Hacks Blocked 15 4 6 3 8 9 11 10 18 15 2 5 4 7 6 8 17 5 15 1 2 18 1 Average FP(%) 0.2 0.6 1.8 0.5 2.6 0 0 2.4 2.4 2.6 0 22.3 4.6 7.9 12.6 0.7 0.4 0.8 1.4 1.5 0.7 1.6 0.9
 Table 6. Summarized Results of Invariants Effectiveness Evaluation.
@@
 Results.
-Table 5 provides a comprehensive summary of the effectiveness evaluation for the invariants applied across various contract categories.
-The table lists several key metrics: the number of contracts to which each type of invariant is applied, the number of contracts successfully protected by these invariants, the number of hacks blocked, and the Average False Positive rate.
-Among these, the row # Hacks Blocked stands out as the most crucial metric as it directly measures the capability of each invariant to block financial exploits.
+Table 5 provides a detailed breakdown of the effectiveness evaluation for each invariant applied across the $42$ contracts.
+The table lists the number of contracts on which an invariant is applied, the number of contracts successfully protected by the invariants, the number of hacks blocked, and the false positive rate for each invariant.
+The row # Hacks Blocked stands out as the most crucial metric as it directly measures the capability of each invariant to block exploits.
+The average false positive rate is also an important metric as it quantifies the potential impact on regular users, reflecting the trade-off between security and usability.
+Table 6 provides a comprehensive summary of the effectiveness evaluation for the invariants applied across various contracts.
+The table lists several key metrics: the number of contracts on which an invariant is applied, the number of contracts successfully protected by the invariants, the number of hacks blocked, and the average false positive (FP) rate.
+Among these, the row # Hacks Blocked stands out as the most crucial metric as it directly measures the capability of each invariant to block exploits.
 The average FP rate is also an important metric as it quantifies the potential impact on regular users, reflecting the trade-off between security and usability.
@@
 These invariants block the highest number of hacks in their respective categories.
-For false positives, EOA, OB and TOU have an average FP rate below 0.4%, while GS and DFU also have a low FP rate below 2.4%.
+For false positives, EOA, OB and TOU have an average FP rate below 0.4%, while GS and DFU have a low FP rate below 2.4%.
 This low rate indicates that these invariants have a small impact on regular user transactions, thereby making them practical for real-world deployment.
@@
 Using upper-bound or range-based invariants for these categories could inadvertently block all transactions once these values exceed a certain threshold.
-Answer to RQ1: EOA in Access Control, OB in Time Lock, GS in Gas Control, TOU in Money Flow, and DFU in Data Flow shows the highest efficacy in blocking hacks while having a low false positive rate.
+Answer to RQ1: EOA in Access Control, OB in Time Lock, GS in Gas Control, TOU in Money Flow, and DFU in Data Flow are the most effective in blocking hacks with low false positive rates.
 6.2. RQ2: Study of False Positives and True Positives
@@
 For each case, we assign one of three categories: C1: the exploit is entirely blocked, and the hacker can no longer gain any profit; C2: the exploit is partially blocked, resulting in significantly reduced profits for the hacker; and C3: the hacker can still achieve profits similar to historical data, with some adjustments to their exploit code.
-We also manually analyze the false positives generated by our invariants to assess their impact on regular, benign users.
-For this, we sample up to 10 transactions from the false positives for each invariant and manually evaluate their bypassability.
-We operate under the assumption that regular users have the option to split transactions with large parameters into transactions with smaller parameters (via DApp’s front-end), lower the gas for their transactions (via a wallet software), or simply wait for some time to transact again after their transaction previously blocked.
-Based on these criteria, we categorize the false positives into three groups: D1: the transaction is completely blocked and cannot be bypassed through simple means; D2: the transaction can be bypassed by breaking it down into smaller transactions; and D3: users can bypass the transaction by reducing the gas or waiting for some time.
-For both true positives and false positives, two authors independently labeled the bypassability results with the third one to resolve the divergence of views.
+We also manually analyze the false positives generated by our invariants to assess their impact on regular users.
+For this, we sample up to 10 transactions from the false positives for each invariant and evaluate their bypassability.
+We operate under the assumption that regular users have the option to split transactions with large parameters into transactions with smaller parameters, lower the gas for their transactions, or simply wait for some time to transact again after their transaction is blocked.
+Based on these criteria, we categorize the false positives into three groups: D1: transaction is completely blocked and cannot be bypassed through simple means; D2: transaction can be bypassed by breaking it down into smaller transactions; and D3: users can bypass the transaction by reducing the gas or waiting for some time.
+For both true and false positives, two authors independently labeled the bypassability results with the third author to resolve divergence of views.
 Table 7. Bypassability Results of Hacks Blocked (TPs) and Sampled Normal Transactions Blocked (FPs).
@@
 OB interestingly has no false positives, but in practice it is very hard to be bypassed by normal users.
-For certain invariants such as TOU and DFU, there is a dichotomy where they block exploit transactions effectively but normal users can still bypass them.
+For certain invariants such as TOU and DFU, there is a dichotomy where they block exploit transactions effectively, but normal users can still bypass them.
 They make certain exploits impossible by preventing hackers from reaching specific contract states required for profitability.
@@
 Building on insights from RQ1 and RQ2, we identify a set of five effective invariants (EOA, SO, GC, TOU, and DFU), replacing GS with GC considering that it can be easily bypassed.
-Our aim is to explore whether these invariants, each effective in its own domain, can be combined to provide a more robust defense against contract exploits while maintaining a low false positive rate.
+Our aim is to explore whether these invariants, each effective in its own domain, can be combined to provide a more robust defense against hacks while maintaining a low FP rate.
 After investigating the exploits blocked by each of these five invariants in RQ1, we have observed that (1) EOA and GC can uniquely block 2 and 1 hacks, respectively; (2) no exploits can uniquely be blocked by OB, TOU or DFU; (3) all exploits blocked by MFU can be blocked by DFU.
@@
 # Hacks Blocked 23 C3 3 # Hacks Blocked 20 C3 2
-Average FP rate(%) 3.99 D1 8 Average FP rate(%) 0.32 D1 10
+Average FP rate(%) 3.99 D1 8 Average FP rate(%) 0.28 D1 10
 D2 2 D2 0
@@
 However, this comes at the expense of a higher false positive rate of 3.99%.
-The combination EOA $\land$ (OB $\lor$ DFU) maintains an impressively low false positive rate of 0.32%, making it superior based on the metric (2).
+The combination EOA $\land$ (OB $\lor$ DFU) maintains an impressively low false positive rate of 0.28%, making it superior based on the metric (2).
 Its efficacy is due in part to OB’s inherent low false positive rate and to the complementary nature of OB and DFU in the context of token transfer functions.
@@
 In those cases, both hackers and normal users may still find a way to bypass the security measures.
-Answer to RQ3: The invariant guards studied and generated by Trace2Inv are complimentary and their combinations are promising to be more effective on contract protection and hack prevention with lower false positive rate.
+Answer to RQ3: The invariant guards studied and generated by Trace2Inv are complimentary and their combinations are promising to be more effective on contract protection and hack prevention with lower false positive rates.
 6.4. RQ4: Gas Overhead of Invariant Guards
 Experiment.
-In our next research question, RQ4, we examine the gas overhead incurred by the deployment of individual and combined invariants.
+In RQ4, we examine the gas overhead incurred by the deployment of individual and combined invariants.
 We select four benchmark contracts that represent different kinds of protocols and programming languages.
@@
 Kind Benchmark Compiler EOA OB GC DFU EOA $\land$ GC $\land$ DFU EOA $\land$ (OB $\lor$ DFU)
-Bridge HarmonyBridge Solidity 0.5.17 0.04 0.52 0.02 0.01 0.07 0.56
-Lending Harvest1 Solidity 0.5.18 0 1.46 0.01 0.01 0.02 1.46
-Yield-Earning CreamFi2_1 Solidity 0.5.17 0 2.55 0 0 0.01 2.55
-Others BeanstalkFarms Vyper 0.2.8 0 0.53 / 0 0 0.53
+Bridge HarmonyBridge Solidity 0.5.17 0.04 0.59 $\rightarrow$ 0.12+ 0.02 0.01 0.07 0.63 $\rightarrow$ 0.16
+Lending Harvest1 Solidity 0.5.18 0.00 0.96 $\rightarrow$ 0.02 0.01 0.01 0.02 0.97 $\rightarrow$ 0.04
+Yield-Earning CreamFi2_1 Solidity 0.5.17 0.00 0.14 $\rightarrow$ 0.01 0.01 0.00 0.01 0.15 $\rightarrow$ 0.01
+Others BeanstalkFarms Vyper 0.2.8 0.00 4.09 $\rightarrow$ 0.68 / 0.00 0.00 4.09 $\rightarrow$ 0.68
 Results.
@@
 This is because the combined invariants do not access new variables, but rather utilize the existing variables to perform additional comparisons.
-Answer to RQ4: The gas overheads of these invariant guards are as low as 0% - 2.55%.
+Answer to RQ4: The gas overheads of these invariant guards are as low as 0% - 0.68%.
 6.5. RQ5: Comparative Analysis with Other State-of-the-art (SOTA) Tools
@@
 Similar to Trace2Inv, InvCon+ takes a target contract and its transaction history as inputs and automatically generates invariants.
-We contacted authors of InvCon+ and obtained the InvCon+ tool.
-We then applied it on our benchmarks in Table 4.
+We obtained InvCon+ (inv, 2024) from its authors, and applied it on our benchmarks in Table 4.
 Following the same methodology of RQ1 and RQ3, we used $70\%$ of the transaction history for training, testing the generated invariants on the remaining $30\%$.
 Experiment 2: Compare with TxSpector, a SOTA transaction attack detection tool.
-Though Trace2Inv is not primarily designed for transaction anomaly detection, we explored its ability to flag attack transactions.
-TxSpector (Zhang et al., 2020), a SOTA framework for transaction attack detection, identifies attacks using eight detectors: re-entrancy, unchecked call, failed send, timestamp dependence, unsecured balance, misuse of origin, suicidal, and gas-related re-entrancy.
-TxSpector takes a transaction as input and automatically flags it if it is identified as an attack.
-We obtained TxSpector from its public repository [^1] https://github.com/OSUSecLab/TxSpector, and applied it to the testing set of our benchmarks.
-Same as what is used in TxSpector paper (Zhang et al., 2020), we assigned a timeout of $60$ seconds for all benign transactions.
-For hack transactions which are usually more complex, we assigned a timeout of $2$ hours.
+Though Trace2Inv is not primarily designed for transaction anomaly detection, we explored its capability to identify attack transactions.
+TxSpector (Zhang et al., 2020) identifies attacks using eight detectors: re-entrancy, unchecked call, failed send, timestamp dependency, unsecured balance, misuse of origin, suicidal, and gas-related re-entrancy.
+We obtained TxSpector from its public repository (txs, 2020).
+TxSpector operates with a modified Geth archive node to produce transaction traces that are fed into the detectors.
+Using components of Trace2Inv along with an unmodified Geth archive node, we generated equivalent transaction traces for the testing sets of our benchmarks, and applied TxSpector detectors to analyze them.
+We also patch TxSpector to support new EVM opcodes introduced after its last update.
+Consistent with the setup in the TxSpector paper (Zhang et al., 2020), we set a timeout of $60$ seconds for all benign transactions.
+We extended this to $2$ hours for hack transactions, which are typically more complex, when applying the detectors.
 Table 10. Comparison among Trace2Inv, InvCon+, and TxSpector. (TxSpector only takes transactions as input, thus it does not have statistics about contracts.)
@@
 # Contracts Protected(42) 31 8 -
-# Hacks(27) 20 Blocked 6 Blocked 7 Detected
+# Hacks(27) 20 Blocked 3 Blocked 7 Detected
 Average # Invariants per Contract 12 2054 -
-Average FP rate(%) 0.32 73.55 15.30
+Average FP rate(%) 0.28 73.55 15.30
 Results.
@@
 Additionally, it inaccurately marked $15.30$% of benign transactions as malicious, compromising its real-world utility.
-In contrast, Trace2Inv, utilizing the invariant template EOA $\land$ (OB $\lor$ DFU), effectively secured 31 out of 42 victim contracts across 20 hacks with a remarkably low false positive rate of just 0.32%.
+In contrast, Trace2Inv, utilizing the invariant template EOA $\land$ (OB $\lor$ DFU), effectively secured $31$ of $42$ victim contracts across $20$ hacks with a remarkably low FP rate of just $0.28$%.
 Moreover, Trace2Inv demonstrated enhanced practicality by generating an average of only $12$ invariants per contract, significantly outperforming InvCon+ in terms of real-world viability.
-Unlike TxSpector, which only detects hacks, Trace2Inv not only blocks a greater number of hacks but also achieves this with a significantly lower false positive rate, indicating its effectiveness in anomaly detection as well.
-Identifying New Exploits: In the development of Trace2Inv, we surprisingly found two previously unreported exploit transactions, earlier than any reported exploit transactions against RariCapital1222https://etherscan.io/tx/0x4764dc6ff19a64fc1b0e57e735661f64d97bc1c44e026317be8765358d0a7392 and Yearn, [^3] https://etherscan.io/tx/0x59faab5a1911618064f1ffa1e4649d85c99cfd9f0d64dcebbc1af7d7630da98b respectively.
-These two transactions were initially reported by Trace2Inv as false positives, because they were not flagged as attacks when we collected benchmarks.
-However, after manual investigation, we found that these two transactions caused a huge financial loss and the addresses of the originators of the two transactions are flagged on EtherScan as “Rari Capital Exploiter” and “Yearn (yDai) Exploiter”, respectively.
-Thus, we believe these are indeed exploit transactions.
-This discovery further underscores Trace2Inv’s potential in unveiling new exploits.
+Unlike TxSpector, which only detects hacks, Trace2Inv not only blocks a greater number of hacks but also achieves this with a significantly lower FP rate, indicating its effectiveness in anomaly detection as well.
+Identifying New Exploits: In the development of Trace2Inv, we surprisingly found two previously unreported exploit transactions, earlier than any reported exploit transactions against RariCapital1 (Rar, 2023a) and Yearn (Yea, 2024).
+The two transactions were initially reported by Trace2Inv as false positives, because they were not flagged as hacks when we collected benchmarks.
+However, after manual investigation, we found the two transactions caused a huge financial loss and the addresses of the originators of the two transactions are flagged on EtherScan as “Rari Capital Exploiter” and “Yearn (yDai) Exploiter”, respectively.
+Thus, we believe they are indeed exploit transactions.
+This discovery underscores Trace2Inv’s potential in unveiling new exploits.
 Answer to RQ5: Trace2Inv outperforms current SOTA works on smart contract invariant mining and transaction attack detection in terms of both practicality and accuracy.
@@
 7. Discussion
-This study contributes valuable insights into the application and effectiveness of invariant-based security measures across different types of decentralized finance (DeFi) protocols.
-In this section, we discuss some key takeaways and their implications for future research and practical applications.
+In this section, we discuss some key takeaways of this work and their implications in the context of invariant-based security measures for decentralized finance (DeFi) protocols.
 Choosing Complementary Invariants.
@@
 Each type of invariant serves as a unique line of defense against abnormal transactions.
-For instance, invariants in time lock category act as temporal barriers, making it difficult for attackers to execute key functions like withdraw multiple times within a single transaction.
-On the other hand, invariants in data flow and money flow categories limit the token amounts that can be withdrawn in one function call.
-By employing a combination of these invariants, developers force attackers to only withdraw a controlled amount of tokens per transaction, which may disrupt the underlying logic that the attack relies on, thereby blocking the attack.
+For instance, invariants in time lock category act as temporal barriers, making it difficult for attackers to execute key functions like withdraw multiple times within one transaction.
+On the other hand, invariants in data and money flows categories limit the token amounts that can be withdrawn in one function call.
+By employing a combination of these invariants, developers force attackers to only withdraw a controlled amount of tokens per transaction, which may disrupt the mechanism the attack relies on, thereby blocking the attack.
 Dynamic Parameter Updates for Invariants.
-Another key insight from our research is the need for dynamic parameter updates for certain types of invariants.
-For invariants that are tied to variables that change cumulatively over time—like oracle prices or storage values—parameters in the invariants can quickly become obsolete.
-If such an invariant is violated, it could lead to a cascade of failed transactions, causing a high False Positive rate.
-Therefore, it is crucial for developers to continuously monitor these variables and adjust the invariant parameters.
+Another key insight is the need for dynamic parameter updates for certain types of invariants.
+For invariants that are tied to variables that change over time—like oracle prices or storage values—parameters can quickly become obsolete.
+If such an invariant is violated, it could lead to a cascade of failed transactions, causing a high FP rate.
+Thus, it is crucial for developers to monitor these variables and adjust the invariant parameters.
 Conversely, for invariants related to independent actions like token transfers, the parameters can remain relatively stable, as user behavior in these domains tends to be stable over time.
 Mitigating Flash Loan Attacks.
-Our benchmarks indicate a significant prevalence of flash loan-based exploits, with 17 out of 27 examined exploits leveraging flash loan.
-Flash loans enable users to borrow large amounts of tokens for the duration of a single transaction, providing attackers with substantial resources to execute complex exploits.
-Our approach of enforcing invariants can effectively mitigate the risk posed by flash loans.
-Many effective invariants such as EOA and OB block flash loan attacks by enforcing attackers to split their transaction into multiple ones.
-Without flash loan, the attacker would need to use their own assets to execute the attack with the uncertainty of other bots’ backrunning between the hackers’ transactions.
+Our benchmarks indicate a significant prevalence of flash loan-based hacks, with $17$ of $27$ examined hacks leveraging flash loan.
+Flash loans enable users to borrow large amounts of tokens for the duration of one transaction, providing attackers with substantial resources to execute complex hacks.
+Our approach can effectively mitigate the risk posed by flash loans.
+Invariants like EOA and OB block flash loan hacks by enforcing attackers to split their transaction into multiple ones.
+Without flash loan, the attacker would need to use their own assets to execute the hack with the uncertainty of other bots’ back-running between the hackers’ transactions.
 This raise not only the technical but also financial barriers to successful attacks.
-The Impact of Invariants on Contract Composability.
-Incorporating invariant guards into smart contracts might limit their adaptability and integration with DeFi protocols.
-However, our study in Section 4 reveals that many invariants stem from existing DeFi protocol requirements, underscoring the preference of developers for security benefits over flexibility.
-Moreover, our findings in RQ2, as discussed in Section 6, show that certain invariants, like OB, has almost no false positives, while others, such as TOU and DFU, can possibly block malicious transactions while allowing normal users to circumvent them.
-These insights imply that with careful selection and understanding of target user behaviors, developers can devise invariant guards that minimally affect contract composability.
+Impact of Invariants on Contract Composability.
+Incorporating invariant guards into smart contracts may limit their adaptability and integration with other DeFi protocols.
+However, our study in Section 4 reveals that many invariants stem from existing DeFi protocols requirements, underscoring the preference of developers for security benefits over flexibility.
+Moreover, our findings in RQ2, as discussed in Section 6, show that certain invariants, e.g., OB, have almost no FP, while others, e.g., TOU and DFU, can possibly block malicious transactions while allowing normal users to circumvent them.
+These insights imply that with careful selection and understanding of user behaviors, developers can devise invariant guards that minimally affect contract composability.
 8. Related Work
 Smart Contract Invariants.
-There exist several works studying smart contract invariants (Liu et al., 2022a, b; Deng et al., 2024; Zhou et al., 2020).
+Prior works have studied various smart contract invariants.
 Zhou et al.
-(Zhou et al., 2020) introduces $6$ invariants to defend against different hacks.
-Cider (Liu et al., 2022a) leverages deep reinforcement learning on smart contract source code to learn invariants that prevent arithmetic overflows.
-SPCon (Liu et al., 2022b) utilizes function callers in past transactions of a contract to recover a likely access control model.
-More recently, Over (Deng et al., 2024) infers safety constraints on oracles from the contract’s source code and the history of oracle updates.
-Compared to the above works, Trace2Inv studies a broader range of invariants and evaluates their effectiveness and bypassability against real-world attacks.
+(Zhou et al., 2020) introduces six invariants to defend against different hacks.
+Cider (Liu et al., 2022a) uses deep reinforcement learning to learn invariants that prevent arithmetic overflows, while SPCon (Liu et al., 2022b) recovers likely access control models from function callers in past transactions.
+Over (Deng et al., 2024) infers safety constraints on oracles using contract source code and oracle update history.
+In contrast, Trace2Inv explores a broader range of invariants and assesses their effectiveness and bypassability against real-world attacks.
 Smart Contract Security Analysis.
-There are a large body of works on detection of smart contract security vulnerabilities (Feist et al., 2019; Brent et al., 2020; Tikhomirov et al., 2018; Sofware Reliability Lab, 2019; Feng et al., 2019; Luu et al., 2016; man, 2019; myt, 2019; Kalra et al., 2018; Jiang et al., 2018; Wang et al., 2020; Wüstholz and Christakis, 2020; Trail of Bits, 2019).
-Oyente (Luu et al., 2016) is one of the earliest symbolic execution-based security tools to detect reentrancy, mishandled exception, transaction order dependence, and timestamp dependence.
+There is a large body of works on the detection of security vulnerabilities in smart contracts.
+Oyente (Luu et al., 2016) is one of the first symbolic execution-based security tools to detect reentrancy, mishandled exception, transaction order dependency, and timestamp dependency.
 It has been extended to detect greedy, prodigal and suicidal contracts (Nikolić et al., 2018).
-Other well-known symbolic-execution tools also include Manticore (man, 2019) and Mythril (myt, 2019) which are able to find other types of vulnerabilities, e.g., dangerous delegatecall, integer overflow, etc.
+Other well-known symbolic-execution tools include Manticore (man, 2019) and Mythril (myt, 2019) which are able to find other types of vulnerabilities, e.g., dangerous delegatecall, integer overflow, etc.
 Slither (sli, 2021) is another popular static security analysis tool for smart contracts.
-It performs data flow and control flow dependency analysis to support up to 87 bug detectors including dangerous strict equality and the dangerous usage of tx.origin.
+It performs data flow and control flow dependency analysis to support up to $87$ bug detectors.
 Other static analyzers include SmartCheck (Tikhomirov et al., 2018) mainly targeting bad coding practices, Securify (Sofware Reliability Lab, 2019) and Ethainter (Brent et al., 2020) for finding information-flow vulnerabilities.
-Moreover, many dynamic analysis tools (Jiang et al., 2018; Atzei et al., 2017; Nguyen et al., 2020; Trail of Bits, 2019; Wüstholz and Christakis, 2020; Wang et al., 2020; Liu et al., 2020) were proposed to detect smart contract vulnerabilities through fuzzing and model-based testing.
-There is a large body of work on the detection and exploitation of DeFi vulnerabilities (Wu et al., 2021; Kong et al., 2023; Qin et al., 2021; Baum et al., 2022; Xue et al., 2022; Deng et al., 2023; Gudgeon et al., 2020).
+Moreover, dynamic analysis tools (Jiang et al., 2018; Atzei et al., 2017; Nguyen et al., 2020; Trail of Bits, 2019; Wüstholz and Christakis, 2020; Wang et al., 2020; Liu et al., 2020) were proposed to detect smart contract vulnerabilities through fuzzing and model-based testing.
+There is a substantial body of work focusing on the detection and exploitation of DeFi vulnerabilities.
 Qin et al.
 (Qin et al., 2021) and Zhou et al.
-(Zhou et al., 2021) formulated oracle price manipulation and arbitrage as an optimization problem to identify attack opportunities maximizing attacker’s profit.
+(Zhou et al., 2021) manually formulated vulnerabilities such as oracle price manipulation and arbitrage into an optimization problem.
+FlashSyn (Chen et al., 2022) proposed a framework that automatically formulates flash loan attacks as a synthesis problem, by approximating the functions of DeFi protocols.
 Wu et al.
@@
 Also, Gudgeon et al.
-(Gudgeon et al., 2020) showcased how to explore flashloan to conduct governance attack.
+(Gudgeon et al., 2020) showcased how to use flashloan to conduct governance attack.
 Baum et al.
-(Baum et al., 2022) surveyed the state-of-the-art mitigation techniques for front-running in DeFi, such as fair ordering, batching of blind inputs, private and secrete state.
+(Baum et al., 2022) surveyed the state-of-the-art mitigation techniques for front-running in DeFi.
 Interestingly, several works (Xue et al., 2022; Deng et al., 2023; Zhang et al., 2023) explored front-running as a defense mechanism against smart contract exploits.
@@
 Runtime Verification and Validation.
-Runtime verification is a light-weight verification approach that monitors the operation of a running system, detects and possibly reacts to unsatisfied observed behaviors against certain properties.
-Runtime verification can be used for validation purpose where the runtime checks are on properties from users’ expectation rather than from formal program specifications (Magazzeni et al., 2017).
+Runtime verification has been used for validation purpose where the runtime checks are on properties from users’ expectation rather than from formal program specifications (Magazzeni et al., 2017).
 Sereum (Rodler et al., 2018) is a general runtime validation framework to protect deployed contracts against reentrancy attacks.
-Sereum extends Ethereum by introducing the detection module for monitoring attacks rooted in different types of reentrancy.
-Solythesis (Li et al., 2020) provides a source-to-source compiler that facilitates the runtime validation of smart contracts.
-Specifically, Solythesis takes as inputs a smart contract code and a user specified invariant.
-It generates an enhanced smart contracts that reject all unexpected contract transactions.
-Additionally, the overhead of the runtime validation approach is shown to be quite small.
-Trace2Inv distinguishes itself from the above-mentioned approaches in that the existing approaches only offer limited protection for several fixed types of attacks.
-In contrast, Trace2Inv is not bound to specific attack types and has been proven effective in mitigating many sophisticated attack vectors.
+Solythesis (Li et al., 2020) provides a source-to-source compiler that compiles smart contracts with user-specified invariants to reject unexpected transactions (Li et al., 2020).
+In contrast, Trace2Inv is not limited to predefined attack types and has demonstrated effectiveness in mitigating a wide range of sophisticated attacks.
 9. Conclusion
 In this paper, we present the first comprehensive study of the effectiveness of practical invariant guards on preventing DeFi smart contract attacks.
-Our large-scale experiments on real-world DeFi hacks demonstrate that the inferred invariant guards are very effective in stopping the existing hacks, but some of the invariant guards can be bypassable by experienced attackers.
-We also found combining multiple invariants can be more effective than individual invariants with a lower false positive rate.
+Our large-scale experiments on real-world DeFi hacks demonstrate that the inferred invariant guards are very effective in stopping the existing hacks, but some invariant guards can be bypassed by experienced attackers.
+We also show combining multiple invariants can be more effective than individual invariants with a lower false positive rate.
 10. Data Availability
-The experimental data of this paper is available on our website (authors of this paper, 2023). The artifact will be open sourced after the paper is accepted.
+All benchmarks, source code, and study statistics are available at our repositories (Chen et al., 2024c, a, d, b). Additional materials of this paper are available in the website (Chen et al., 2024e).
+Acknowledgement
+This work was supported by MITACS Accelerate program, NSERC Postdoctoral Fellowship, Singapore Ministry of Education Academic Research Fund Tier 1 (RG12/23), and Nanyang Technological University Centre in Computational Technologies for Finance (NTU-CCTF).